### v2

# DBCopilot: Scaling Natural Language Querying to Massive Databases

Abstract.
NL2SQL simplifies database interactions by enabling non-experts to convert natural language (NL) questions into Structured Query Language (SQL) queries. While recent advances in large language models (LLMs) have improved the zero/few-shot NL2SQL paradigm, existing methods face scalability challenges when dealing with massive databases. This paper introduces DBCopilot, a framework that addresses these challenges by employing a compact and flexible copilot model for routing over massive databases. Specifically, DBCopilot decouples schema-agnostic NL2SQL into domain-specific schema routing and generic SQL generation. This framework utilizes a lightweight differentiable search index to construct semantic mappings for massive database schemas and navigates natural language questions to their target databases and tables in a relation-aware, end-to-end manner. The routed schemas and questions are then fed into LLMs for effective SQL generation. Furthermore, DBCopilot introduces a reverse schema-to-question generation paradigm that can automatically learn and adapt the router over massive databases without manual intervention. Experimental results demonstrate that DBCopilot is a scalable and effective solution for schema-agnostic NL2SQL, providing a significant advance in handling large-scale schemas of real-world scenarios.
Text-to-SQL, LLM, Generative Retrieval, Schema Routing
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXccs: Information systems Structured Query Languageccs: Information systems Top-k retrieval in databasesccs: Information systems Question answering
Figure 1. The DBCopilot framework. DBCopilot decouples NL querying over massive databases into schema routing and SQL generation through LLM-Copilot collaboration.
1. Introduction
Natural language to SQL (NL2SQL) aims to convert natural language (NL) questions into Structured Query Language (SQL) queries, making it convenient for non-experts to query and analyze databases. This task plays a critical role in data analysis, business intelligence, and other database interaction applications, thus has attracted significant attention (Androutsopoulos et al., 1995; Affolter et al., 2019; Qin et al., 2022; Katsogiannis-Meimarakis and Koutrika, 2023; Fu et al., 2023; Gu et al., 2023a). The recent emergence of large language models (LLMs) brings a new zero/few-shot NL2SQL paradigm that directly transforms SQL queries via LLMs with schema-aware prompts, showing great promise (Rajkumar et al., 2022; Liu et al., 2023a; Pourreza and Rafiei, 2023; Gu et al., 2023b; Gao et al., 2023).
However, existing work in this direction typically focuses only on querying a single database with a handful of tables, neglecting the real-world requirement to query over massive databases and data warehouses (Pride, 2023; Kothyari et al., 2023; Sen et al., 2020). In this case, data consumers need to invest significant effort to manually identify and select the appropriate database and relevant tables. Therefore, it becomes imperative to develop a mechanism that abstracts non-expert users from the intricacies of data representation and enables them to query over massive databases with schema-agnostic NL questions, i.e., without understanding of large-scale schemas. The core of this process is to construct semantic mappings from diverse NL questions to massive database model elements, thus automatically identifying the target database and filtering for a minimal amount of relevant tables.
The above scenarios pose significant challenges to current LLM-based NL2SQL methods. Due to token limits and inference costs, it is infeasible to feed hundreds/thousands of schemas into LLMs for each question. Furthermore, current LLMs still struggle to effectively leverage information present in long contexts (Liu et al., 2023b; Xu et al., 2023). To address this scalability problem, existing solutions can be categorized into retrieval-based and fine-tuning based methods. Given an NL question, retrieval-based methods identify the corresponding table schemas using information retrieval techniques (Wikipedia contributors, 2023a; Karpukhin et al., 2020). Unfortunately, by relying on the co-occurrence of words and representations between question and schema, retrieval-based methods suffer from vocabulary mismatch and the diversity of NL expressions. In addition, these ad hoc retrieval methods typically treat documents as independent objects, neglecting the relationships between databases and tables. Fine-tuning based methods inject schema knowledge into the parameters of LLMs by fine-tuning with schema descriptions or (question, schema) pairs. Unfortunately, fine-tuning LLMs with billions of parameters to keep up with evolving database schemas is resource-intensive and sometimes impractical. Besides, it remains challenging to effectively inject new knowledge into LLMs through fine-tuning (Ovadia et al., 2023). As a result, how to extend the scalability of LLM-based NL2SQL methods and apply them to massive databases is a major challenge in this field.
In this paper, we present DBCopilot, a framework for effectively scaling natural language querying to massive databases. The main idea behind DBCopilot is to introduce a compact and flexible copilot model to serve as a router across massive databases, and to leverage the efficient NL2SQL capability of LLMs to generate final SQL queries. To this end, as shown in Figure 1, DBCopilot decouples schema-agnostic NL2SQL into domain-specific schema routing and generic SQL generation. Specifically, DBCopilot utilizes a lightweight differentiable search index (DSI) (Tay et al., 2022) to navigate an NL language question to its corresponding database and tables. Then the routed schemas and questions are fed into LLMs to generate SQL queries by leveraging their strong language understanding and SQL generation capabilities. In this way, DBCopilot can easily integrate with advanced LLM-based NL2SQL solutions and rapidly and cost-effectively scale to massive databases.
For an effective implementation of DBCopilot, three major challenges need to be addressed. (1) Schema routing is not a task of retrieving independent tables because complex queries require information across multiple tables, and there are inherent relationships between databases and tables. To resolve this challenge, we model schema routing as a sequence-to-sequence (Seq2Seq) DSI. By proposing schema graph construction, DFS serialization, and constrained decoding, we can jointly retrieve target databases and tables in an end-to-end manner, taking into account structural relationships. (2) The generative retrieval paradigm makes it infeasible to perform semantic mapping of unseen schemas because they are not indexed into model parameters. In practice, labeling NL questions for massive database schemas is highly expensive. To this end, we propose a reverse schema-to-question generation paradigm to automatically synthesize training data and efficiently train the schema router. (3) How to incorporate routed schemas with LLMs to perform schema-aware SQL generation is critical to the quality of NL2SQL. We explore various prompt strategies to select and incorporate multiple candidate schemas for DBCopilot.
We use three standard NL2SQL datasets and two additional robustness datasets to evaluate DBCopilot. Experimental results verify the effectiveness of DBCopilot, with its copilot model outperforming retrieval-based baselines in schema routing by up to 19.88% in recall, and its schema-agnostic NL2SQL showing an improvement in execution accuracy of more than 7.35%. In summary, we demonstrate that DBCopilot is a scalable and effective solution for schema-agnostic NL2SQL, presenting a significant step forward in handling large-scale schemas of real-world scenarios.
Contributions.  Our contributions can be summarized as follows.
  • We highlight the limitations of existing NL2SQL methods in scaling from limited tables to massive databases.
  • We present DBCopilot, a framework that decouples schema-agnostic NL2SQL into schema routing and SQL generation via an LLM-copilot collaboration architecture.
  • We propose a relation-aware, end-to-end joint retrieval approach for schema routing, and also propose a reverse generation paradigm for automatic training data synthesis.
  • We conduct thorough experiments to verify the effectiveness of our solution in scaling NL2SQL to massive databases.
2. Preliminaries
2.1. Problem Definition
Table 1. Notation list used in this paper.
Symbol Description
$N$ Natural language question
$Q$ SQL query equivalent to question $N$
$\mathcal{D}$ Set of databases available, each denoted by $D$
$T$ Set of tables involved in query $Q$, with their properties
$S$ SQL query schema $\langle D,T\rangle$, where $T$ is the subset of tables belonging to database $D$
$\mathcal{T}$ Union of sets of tables from each database in $\mathcal{D}$
$\mathcal{S}$ Schema space $\mathcal{D}\times\mathcal{P}(\mathcal{T})$, where $\mathcal{P}(\mathcal{T})$ is the power set of $\mathcal{T}$
$\mathcal{G}$ Schema graph $\langle\mathcal{V},\mathcal{E}\rangle$, where $\mathcal{V},\mathcal{E}$ are graph nodes and edges
As mentioned above, we formulate the problem of natural language querying over massive databases and large numbers of tables with unknown SQL query schema (i.e., target database and tables) as schema-agnostic NL2SQL. Formally, as per the notation list detailed in Table 1, the problem of schema-agnostic NL2SQL is defined in Definition 2.1, and its two subtasks schema routing and SQL generation are defined in Definition 2.2 and Definition 2.3, respectively. Obviously, traditional NL2SQL with an NL question and its limited database schema as input is similar to the subtask SQL generation.
Definition 0 (Schema-Agnostic NL2SQL).
Given a natural language question $N$, schema-agnostic NL2SQL generates an equivalent SQL query $Q$ that can be executed against the appropriate database $D$ drawn from a collection of databases $\mathcal{D}$, without being provided with the intended SQL query schema $S=\langle D,T\rangle$.
Definition 0 (Schema Routing).
Schema routing is the process of identifying the SQL query schema $S=\langle D,T\rangle$ of a natural language question $N$ from the entire space of all possible schemas $\mathcal{S}=\mathcal{D}\times\mathcal{P}(\mathcal{T})$ within the specified set of databases $\mathcal{D}$ and tables $\mathcal{T}$.
Definition 0 (SQL Generation).
SQL generation is the process of generating a SQL query $Q$, given a natural language question $N$ and the intended SQL query schema $S=\langle D,T\rangle$.
Schema Linking vs. Schema Routing.  Schema linking (Katsogiannis-Meimarakis and Koutrika, 2023) is an optional substep of traditional NL2SQL methods. Given an NL question and a database schema, schema linking aligns NL terminologies, called “mentions”, with their corresponding schema elements (such as tables, columns, or database values). In this way, schema linking servers as a bridge between NL questions and database elements.
Schema linking and schema routing can be distinguished in two aspects. First, schema linking is a substep of SQL generation for accurate generation, while schema routing is a prerequisite of SQL generation to identify the unknown schema of a question. Second, the input for schema linking is the NL question along with the schema, whereas, the input for schema routing is the NL question and the output is its indexed or memorized schema.
2.2. Generative Retrieval
Traditional information retrieval (IR) techniques typically follow the “index-retrieve” paradigm. However, such a pipelined framework, which cannot be optimized together in an end-to-end manner, limits the final retrieval quality. To address this concern, generative retrieval (Tay et al., 2022; Wang et al., 2022) (a.k.a. DSI) is an emerging paradigm that trains a Seq2Seq pre-trained language model (PLM) to map queries directly to relevant document identifiers. The main idea behind it is to completely parameterize traditional “index-retrieve” pipelines within a single neural model, by modeling indexing and retrieval operations into the model parameters. Critical ongoing research in this paradigm focuses on defining better document identifiers (Sun et al., 2023; Zhang et al., 2023). To the best of our knowledge, this paper is the first to adapt this paradigm to jointly retrieve multiple elements with structural relationships.
3. Methodology
This section introduces DBCopilot, a schema-agnostic NL2SQL framework for effectively scaling NL querying to massive databases.
3.1. Overview
As shown in Figure 1, DBCopilot decouples natural language querying over massive databases into schema routing and SQL generation through LLM-Copilot collaboration. Specifically, given an NL question $N$, DBCopilot utilizes a copilot model as a schema router to identify the SQL query schema $S=\langle D,T\rangle$ associated with that question. After that, DBCopilot utilizes an LLM to translate the NL question $N$ with the routed schema $S$ into a corresponding SQL query $Q$. Example 3.1 provides an illustration of this process.
Example 0.
Given a question “Which singers held concerts in 2022?”, DBCopilot converts it into an equivalent SQL query through the following two phases.
1. Schema Routing identifies the relevant database and tables that contain the information required to answer the question:
[…]
2. SQL Generation prompts the LLM with the identified schema and question, and generates the SQL query:
SELECT s.name FROM singer_in_concert AS sic JOIN singer AS s
ON sic.singer_id = s.singer_id JOIN concert AS c
ON sic.concert_id = c.concert_id WHERE c.year = 2022
Figure 2. The training process of the schema router.
In practice, the schema space $\mathcal{S}=\mathcal{D}\times\mathcal{P}(\mathcal{T})$ is huge because it is the cartesian product of the database collections $\mathcal{D}$ and the power set of all tables $\mathcal{T}$. To narrow down the search space, we propose a relation-aware, end-to-end joint retrieval approach for schema routing. The training process is illustrated in Figure 2. Specifically, we first construct a schema graph to represent the relationships between databases and tables (§ 3.2). Based on this graph, we design a schema serialization algorithm to serialize a SQL query schema into a token sequence so that it can be generated by Seq2Seq DSI (§ 3.3). To address the challenge that DSI cannot generalize to unseen schemas, we further propose a reverse generation paradigm for training data synthesis (§ 3.4). These measures allow us to effectively train the schema router using serialized pseudo-instances ¡$N$, $S$¿ and apply inference with graph-based constrained decoding (§ 3.5). Finally, we explore various prompt strategies to select and incorporate multiple candidate schemas with LLMs (§ 3.6).
3.2. Schema Graph Construction
Unlike NL expressions, a valid SQL query must follow the structure of the database. Therefore, it is necessary to inform the schema router about the inherent relationships/constraints between schema elements. In general, the schema $S=\langle D,T\rangle$ involved in a valid SQL query must satisfy the following conditions. First, the set of tables $T$ must be the subset of tables belonging to database $D$, since a SQL query obviously cannot reference tables that do not exist in the target database. Second, the tables in $T$ should cross-reference at least one other table with a common column (e.g., foreign key), so that the SQL query can link them together via join or subquery clauses. In addition to explicit Primary-Foreign relations, two tables can also be linked by implicit Foreign-Foreign relations, i.e., they reference the same unique key column of another table.
Example 0.
Given a question “Which rivers run through the state with the largest city in the US?”, the corresponding SQL query “SELECT river_name FROM river WHERE traverse IN ( SELECT state_name FROM city WHERE population = ( SELECT MAX ( population ) FROM city ) )” uses only the city and river tables, both of which refer to the primary key of the state table.
To this end, we construct a directed graph $\mathcal{G}=\langle\mathcal{V},\mathcal{E}\rangle$ to represent the underlying schema structure of all databases $\mathcal{D}$ and their tables $\mathcal{T}$. Concretely, the schema graph consists of three types of nodes $\mathcal{V}=\{\nu_{s}\}\cup\mathcal{D}\cup\mathcal{T}$, where $\nu_{s}$ is a special node denoting the set of all databases $\mathcal{D}$. Each edge $e\in\mathcal{E}$ in this graph represents a relation between two nodes, which can be divided into:
  • Inclusion relations: indicate whether a database is part of a database collection and whether a table belongs to a database.
  • Table relations: include explicit Primary-Foreign and implicit Foreign-Foreign relations between table nodes.
1
Input: A dictionary of all database schemas $databases$.
Output: The heterogeneous directed schema graph $\mathcal{G}$.
2 $\mathcal{G}\leftarrow$ an empty directed graph
3 for $database$ in $databases$ :
4       $links\leftarrow$ a dictionary with empty set as default value
5       $\mathcal{G}$.addEdge($\nu_{s}$, $database$)
6       for $table$ in $database$ :
7             $\mathcal{G}$.addEdge($database$, $table$)
8             for $column$ in $table$ :
9                   if $column$ contains a foreign key :
10                         $ltbl$, $lcol$leftarrow $column.{\mkern-2.0muf\mkern-2.0mu}oreign\_key$
11                         Add $ltbl$ and $table$ into set $links[ltbl.lcol]$
12
13
14
15
16
17
18      for $tables$ in $links$ :
19             for $tbl1$, $tbl2$ in $tables\times tables$ :
20                   if $tbl1$ is not equal to $tbl2$ :
21                         $\mathcal{G}$.addEdge($tbl1$, $tbl2$)
22
23
24
25
26
27
28
29
30return $\mathcal{G}$
Algorithm 1 Schema Graph Construction.
Algorithm 1 shows the procedure for constructing this three-tiered hierarchical schema graph. By capturing the structure of database schemas in graph $\mathcal{G}$, any valid SQL query schema $S$ is a trail (Wikipedia contributors, 2023b) on this graph. Each schema trail starts with node $\nu_{s}$, followed by a database node and some of its interconnected table nodes. The schema graph can efficiently narrow down the schema space $\mathcal{S}$ into trials on this graph starting from node $\nu_{s}$. As a result, the schema graph is essential for effectively representing and sampling SQL query schemas, serving as the foundation of DBCopilot.
3.3. Schema Serialization
Since we model schema routing as a Seq2Seq generative retrieval process, we need to design a schema serialization algorithm that converts a SQL query schema into a sequence of tokens.
Basic Serialization.  A trivial solution is to randomly order the set of tables as ad hoc retrieval (Chen et al., 2023), and identify the database based on the votes of the generated tables. However, this method loses sight of the inclusion and table relations between schema elements. The former results in a huge search space due to the lack of database constraints during table generation. The latter can result in missing tables during schema routing, especially junction tables.
Depth-First Search Serialization.  To address the above problem, we follow the constructed schema graph and design a depth-first search (DFS) serialization algorithm. As shown in Algorithm 2, we perform a DFS starting at node $\nu_{s}$ until all nodes of a SQL query schema have been visited and record the access order. We then concatenate the sequence of node names into a sequence of tokens. Formally, the depth-first search serialization can be defined as:
$\operatorname{Serialize}_{\pi}(S)\coloneqq\operatorname{Concat}\left(\left[\operatorname{DFS_{\pi}}(S,\mathcal{G})[i]\right]_{i=1}^{\left|\operatorname{DFS_{\pi}}(S,\mathcal{G})\right|}\right)$
where $\pi$ is a node iteration order and $\operatorname{DFS_{\pi}}(S,\mathcal{G})$ is the DFS traversal sequence of schema $S$ on schema graph $\mathcal{G}$ when iterating successor nodes under order $\pi$.
We introduce the iteration order parameter $\pi$ to determine the DFS sequence because a SQL query schema is not a completely ordered set and there may be multiple DFS sequences for a schema. We randomly select one DFS sequence at a time as the model target during training and assign different probabilities to different sequences originating from the same schema during inference.
1
Input: An ordered pair of a SQL query schema $S=\langle D,T\rangle$, the constructed schema graph $\mathcal{G}$, and a node iteration order $\pi$.
Output: The serialization of schema $S$.
2 $nodes$leftarrow $\{\nu_{s}\}\cup\{D\}\cup T$
$visited$leftarrow [][\;][ ]
  /* an empty list */
3 $stack$leftarrow $[\nu_{s}]$
4 while $stack$ is non-empty :
5       $node$leftarrow $stack$.pop()
6       $visited$.append($node$)
7       if set($visited$) is equal to $nodes$ :
             break
              /* cease the loop */
8
9
10      $successors$leftarrow [][\;][ ]
11       for $successor$ in $\mathcal{G}$[$node$] :
             /* Iterates over successor nodes according to node iteration order $\pi$/
12             if $successor$ in $nodes$ and not in $visited$ :
13                   $successors$.append($successor$)
14
15
16
17
18      $stack$.extend($successors$)
19
20
return $\operatorname{Concat}(visited[1:\operatorname{len}(visited)])$
  /* skip $\nu_{s}$/
Algorithm 2 Depth-First Search Serialization.
3.4. Training Data Synthesis
Based on the schema graph and the DFS serialization, we can learn the schema router on question-schema pairs ($N$, $S$) as described in § 3.5. However, the generative retrieval paradigm prevents the schema router from generalizing to unseen schemas because they are not injected into the model parameters. To address this challenge, we propose a labor-free training data synthesis method that automatically generates pseudo-instances via a reverse schema-to-question generation paradigm, as shown in Figure 2.
Specifically, we sample a sufficient number of valid schemas from the schema graph and generate a pseudo-question for each sampled schema. These schemas are sampled by performing finite-length random walks over the schema graph, and the database and tables traversed by a walk form a sampled schema. To generate an NL question for each sampled schema, we train a schema questioning model $\mathcal{M}_{\text{q}}$ in reverse based on the training sets of NL2SQL datasets. We use a SQL parser to extract the metadata (tables and columns) of SQL queries from the training pairs ($N$, $Q$). Unlike the schema router, which takes a condensed SQL query schema as output, the schema questioner accepts more detailed schema information (such as table columns) as input to generate rich and detailed questions. Given an illustration of the model input and output in Figure 3, the training loss of the schema questioning model $\mathcal{M}_{q}$ with parameters $\theta_{\text{q}}$ on $m$ extracted schema-question pairs ($S$, $N$) is:
$\mathcal{L}_{q}=-\sum_{i=1}^{m}\log P\left(N_{i}\mid S_{i};\theta_{\text{q}}\right)$
Pseudo-Question Generation Input: Ask a question that needs to be answered by combining the contents of all the tables, based on the word1 database schema provided below. - countrylanguage(language, countrycode, official) - country(code, lifeexpectancy) Output: What is the official language of the country with the highest life expectancy?
Figure 3. An illustration of pseudo-question generation.
3.5. Schema Router Training & Inference
Through the above measures, we can efficiently train the schema router to parameterize the semantic mapping of massive database schemas and effectively route NL questions to their target schemas via graph-based constrained decoding. Formally, the Seq2Seq schema router $\mathcal{M}_{r}$ takes a question sequence $N=\left[t_{1},t_{2},\ldots,t_{|N|}\right]$ as input and generates the serialized schema sequence $\bar{S}=\left[s_{1},s_{2},\ldots,s_{|\mathbf{s}|}\right]$. Suppose the synthetic data is $\{(N_{i},S_{i})\mid 1\leq i\leq n\}$, the training loss of schema router $\mathcal{M}_{r}$ with parameters $\theta_{r}$ is:
$\mathcal{L}_{r}=-\sum_{i}^{n}\sum_{\pi}\log P\left(\bar{S}_{i}\mid N_{i};\theta_{r}\right)$
After training the schema router, we further design a graph-based constrained decoding algorithm to guarantee the validity of generated schemas. As shown in Figure 4, at each step of autoregressive decoding, we maintain a dynamic prefix tree containing the names of accessible nodes from decoded schema elements. Unlike the DFS serialization, the accessible table nodes can only be the neighbors of generated tables, not the neighbors of the database, unless no tables have been generated yet. Therefore, the available tokens (part of the node name) for each decoding step can be obtained by searching the prefix tree, prefixed with the tokens generated after the last element separator. Besides, multiple candidate schema sequences can be generated to improve the coverage of the target schema. We thus utilize diverse beam search (Vijayakumar et al., 2016), which allows for generating a more diverse set of candidate sequences.
Figure 4. An example of graph-based constrained decoding.
3.6. SQL Generation
By decoupling schema-agnostic NL2SQL into schema routing and SQL generation, DBCopilot can easily integrate with orthogonal advanced LLM-based NL2SQL solutions, no matter whether in-context prompt engineering methods (Pourreza and Rafiei, 2023; Gao et al., 2023; Guo et al., 2023; Chang and Fosler-Lussier, 2023) or specific NL2SQL LLMs (Defog.ai, 2023; Station, 2023). However, since the schema router can generate multiple schemas from different databases for an NL question, we explore three prompt strategies to select and incorporate these different database schemas for LLMs:
  • Best Schema Prompting: instructs LLMs with the highest probability schema generated by the schema router.
  • Multiple Schema Prompting: instructs LLMs using multiple candidate schemas by simply concatenating the table schemas from different sequences in the prompt.
  • Multiple Schema COT Prompting: instructs LLMs using multiple candidate schemas through chain-of-thought (COT) reasoning (Wei et al., 2022). The LLM is first asked to select the most relevant candidate schema and then generates SQL queries.
4. Experiments
In this section, we conduct experiments on schema routing and schema-agnostic NL2SQL to verify the effectiveness of DBCopilot.
4.1. Experimental Setup
4.1.1. Datasets
We adapt existing NL2SQL datasets to evaluate schema routing and schema-agnostic NL2SQL.
  • Spider (Yu et al., 2018): Spider is a widely-used cross-domain NL2SQL dataset with 10,181 questions and 5,693 unique SQL queries across 138 domains and 200 databases.
  • Bird (Li et al., 2023b): Bird is a recent challenging NL2SQL dataset designed to emphasize the importance of database values, spotlighting the challenges presented by dirty database content, external knowledge requirements, and SQL efficiency.
  • Fiben (Sen et al., 2020): Fiben is a dataset designed to evaluate the processing of complex business intelligence queries, often involving nesting and aggregation, for NL2SQL systems. It consists of 300 NL questions, corresponding to 237 different complex SQL queries on a database with 152 tables.
  • Spidersyn (Gan et al., 2021): It is a human-curated dataset designed to evaluate the robustness of NL2SQL models to synonym substitution. It builds on Spider with NL questions modified by replacing schema-related words with real-world paraphrases.
  • Spiderreal (Deng et al., 2021): This is another robustness dataset created from the dev split of the Spider dataset by manually modifying the original questions to remove the explicit mention of column names while leaving the SQL queries unchanged.
Table 2. Statistics of datasets. Asterisks (*) indicate that the original development sets are used as test sets.
Dataset Size Schema
Train Dev. Test # DB # Tables # Cols
Spider 7000 1034* 166 876 4503
Bird 9427 1534* 80 597 4337
Fiben - - 279 1 152 374
Spidersyn 7000 1034 166 876 4503
Spiderreal 508 166 876 4503
Dataset Adaptation.  These datasets are designed to evaluate NL2SQL on a single database, where each question is annotated with its target database. To approach the schema-agnostic NL2SQL, we make several adjustments. First, we remove the single-database constraint and treat all databases in the collection as potential query targets. Next, we extract metadata (tables and columns) from each SQL query using a SQL parser (Mao, 2023) and exclude any queries that cannot be parsed. Finally, we combine the target database and extracted metadata as a SQL query schema, together with the original NL question and SQL query, to form each instance ($N$, $S$, $Q$). This results in three database collections (Spider, Bird, and Fiben), with instances divided into training and test sets as in the original setup. The robustness variants of Spider share the same database collection. Table 2 summarizes the dataset statistics after adaptation.
Table 3. Schema routing performance on standard test sets. The best results are in bold and the second best are underlined.
Method Spider Bird Fiben
Database Table Database Table Database Table
R@1 R@5 R@5 R@15 R@1 R@5 R@5 R@15 R@1 R@5 R@5 R@15
Zero-shot
BM25 70.12 91.49 86.49 93.87 57.11 90.16 68.32 82.81 98.92 98.92 33.34 38.62
SXFMR 62.57 89.46 80.42 92.39 66.49 88.20 67.56 83.05 100.0 100.0 28.21 46.47
LLM-enhanced
[…]
CRUSHSXFMR 63.25 91.97 82.21 93.86 70.14 92.44 70.58 85.07 100.0 100.0 34.06 50.78
Fine-tuned
BM25 68.28 91.78 86.60 93.87 60.63 90.48 69.87 83.34 98.92 98.92 32.67 38.30
[…]
4.1.2. Baselines
We compare DBCopilot to various sparse/dense retrieval techniques for scaling NL querying over massive databases, including zero-shot, LLM-enhanced, and fine-tuned methods:
  • BM25 (Wikipedia contributors, 2023a): BM25 is a well-known ranking function widely used in search systems to retrieve relevant documents.
  • SXFMR (Reimers and Gurevych, 2019): SXFMR is a technique for obtaining generic embedding models for dense retrieval by contrastive learning of Transformer(XFMR)-based PLMs on related text pairs.
  • CRUSH (Kothyari et al., 2023): CRUSH adopts a two-stage approach to retrieve schema elements for NL2SQL. It first instructs the LLM to generate a hallucinated schema, and then combines multiple retrieval results to obtain the actual schema.
  • DTR (Herzig et al., 2021): DTR is an approach that uses contrastive learning on labeled (question, table) pairs to develop table retrievers for open-domain question answering over tables.
4.1.3. Evaluation Metrics
For schema routing, we use Recall@k and mAP (mean average precision) to evaluate the relevance of retrieved databases and tables. Recall@k measures the fraction of relevant instances in the top-k retrieved candidates. mAP calculates the average precision over all queries, which reflects the overall retrieval performance. For SQL generation, traditional surface-level metrics are inappropriate for LLM-based approaches as they may produce SQL queries that differ greatly from the ground truth. Following recent studies (Gu et al., 2023b; Gao et al., 2023; Li et al., 2023b), we use EX (execution accuracy), which compares the results of generated SQL queries with corresponding ground truths. We also report the cost of LLM invocations.
4.1.4. Implementation Details
We implemented DBCopilot using Python 3.10, Pytorch 2.1, and the Transformers (Wolf et al., 2020) library.
Schema Routing.  We adopted T5-base (Raffel et al., 2020) as the backbone of the schema questioner and router. We set the batch size to $32$, the learning rate to $5\times 10^{-5}$, and the maximum number of training epochs to $20$ for all models. Model parameters were optimized using AdamW (Loshchilov and Hutter, 2019) and a linear learning rate schedule with no warm-up steps. We used the schema questioner trained on Spider and Bird to generate pseudo-questions for Fiben, which has no training data. For each database collection, we synthesized $10^{5}$ training instances.
SQL Generation.  We used gpt-3.5-turbo-16k-0613 to select the optimal schema and generated the SQL query with the OpenAI demo prompt (Gao et al., 2023). We set the temperature to 0 for reproducibility.
Baselines.  To maintain fairness, all baselines compared in this study relied on schema information only. We treated tables as retrieval targets, with the content of each table being the flat normalized names of the table and its columns. For each NL question, we retrieved the top 100 tables and determined database scores based on the average of their table scores. We used the standard Okapi BM25 with two adjustable parameters for BM25 and adopted the best performing all-mpnet-base-v2 for SXFMR. BM25 and DTR were fine-tuned on synthetic data, consistent with DBCopilot.
4.2. Schema Routing
We first compare the performance of schema routing, with the results shown in Table 3, Table 4, Figure 5, and the following findings.
Table 3 shows the database and table recall on regular test sets. We can see that DBCopilot consistently outperforms all types of baselines by a large margin. The database recall@1 of DBCopilot surpasses the second-best approach by up to 12.67% and 11.93% on the Spider and Bird datasets, respectively. This result verifies that DBCopilot can effectively identify the target database of an NL question. Furthermore, DBCopilot also holds an advantage in table recall, with a lead of up to 9.59%, which is crucial for providing complete tables for SQL generation. Even on the single database dataset Fiben, DBCopilot still achieves a 3.4% improvement. These results confirm that DBCopilot can effectively navigate NL questions to their target databases and tables over massive databases.
Finding 1. DBCopilot consistently outperforms all types of baselines by a large margin in schema routing.
Table 4. Schema routing performance on robustness test sets.
Method Spidersyn Spiderreal
Database Table Database Table
R@1 R@5 R@5 R@15 R@1 R@5 R@5 R@15
Zero-shot
BM25 29.11 58.03 45.30 58.87 52.95 81.1 70.87 80.64
[…]
LLM-enhanced
CRUSHBM25 35.88 65.18 46.80 64.06 58.46 86.42 75.38 89.14
[…]
Fine-tuned
BM25 30.66 57.83 46.12 59.65 50.00 81.30 70.77 80.87
[…]
DBCopilot 62.67 85.11 70.35 86.26 79.53 95.47 87.93 96.06
When querying over massive databases, data consumers without schema knowledge typically express NL questions that are highly inconsistent with database schemas. To evaluate the robustness in this situation, Table 4 shows the results of Spidersyn and Spiderreal, which replace explicitly mentioned schema elements in NL questions posed by NL2SQL annotation. We can see that the performance of all methods decreases with this diversity setting. Sparse retrieval baselines are significantly affected as they rely heavily on the vocabulary matching assumption, and the LLM-enhanced method also struggles with this challenge, despite being designed specifically for this case. Compared to these baselines, DBCopilot is the least affected and therefore maintains a larger advantage with 14.89% and 19.88% on database recall@1, as well as 2.83% and 6.94% on table recall@5. We believe this is because by injecting schema knowledge through schema-specific training, our method can better understand the semantics of database schemas and effectively learn the mapping between questions and schemas.
Finding 2. DBCopilot is more robust when natural language questions are highly inconsistent with database schemas.
Figure 5. (a) Table mAP w.r.t. different numbers of database tables. (b) Table recall@$k$ w.r.t. the number of retrieved tables.
DBCopilot model schema routing as a relation-aware, generative joint retrieval process, which differs from traditional baselines of independent retrieval. To assess the effect of relation modeling and joint retrieval, we evaluate the performance under different numbers of database tables and retrieved tables. As shown in Figure 5 (a), the mAP of DBCopilot is relatively stable as the number of database tables increases, while the independent retrieval baselines decrease significantly. The recall of DBCopilot is also consistently higher than the baselines as the number of retrieved tables increases from Figure 5 (b). These results verify the scalability of DBCopilot through relation modeling and generative routing.
Finding 3. Relation modeling and joint retrieval are essential for DBCopilot to perform effective schema routing.
4.3. Schema-Agnostic NL2SQL
We then compare the schema-agnostic NL2SQL performance under gold schemas, different routing methods, and prompt strategies.
Oracle Test.  We conduct oracle tests to access the upper bound of schema-agnostic NL2SQL, by providing gold schemas at various small scales. Specifically, we gradually move from 5 database schemas to the gold database, then to gold tables only, or even to gold tables with only gold columns. Experimental results in Table 5 show that as the number of schema elements increases, the EX of LLM-generated SQL queries continuously decreases, while the inference cost keeps increasing. This suggests that LLMs, like PLMs (Li et al., 2023c), are also affected by extraneous schema elements in SQL generation. Predictably, the situation gets worse as the scale of the schema increases. As a result, it is highly inefficient and ineffective to expose massive database schemas directly to LLMs.
Table 5. Schema-agnostic NL2SQL performance with gold (G.) schemas, different routing methods, and prompt strategies.
Routing Method Provided Schema Prompt Strategy Spider Bird Spidersyn
EX Cost EX Cost EX Cost
G. Column Basic 79.69 0.40 35.98 0.75 76.60 0.40
G. Table Basic 77.08 0.52 29.86 1.44 67.21 0.52
G. Database Basic 74.47 0.81 25.75 2.70 63.73 0.81
5 Database Basic 67.89 2.98 19.10 8.80 49.90 2.99
CRUSHBM25 Top 1 Basic 56.77 1.30 13.69 2.67 20.79 1.23
DTR Top 1 Basic 45.16 0.66 17.60 2.12 26.98 0.62
DBCopilot Top 1 Basic 64.12 0.69 21.97 2.35 37.81 0.70
DBCopilot Top 5 Basic 61.51 2.28 19.56 7.18 34.53 2.27
DBCopilot Top 5 COT 57.83 5.33 19.17 16.49 33.95 5.29
[…]
Finding 4. Schema routing is necessary for LLM-based schema-agnostic NL2SQL for both efficiency and effectiveness.
Impact of Improved Recall.  We perform schema-agnostic NL2SQL to verify the impact of improved recall on the accuracy of generated SQL queries. Without loss of generality, we compare DBCopilot to CRUSHBM25 (sparse) and DTR (dense), which achieve suboptimal routing performance on Spider and Bird, respectively. As shown in Table 5, DBCopilot outperforms these two baselines by up to 10.83% EX. The LLM inference cost of DBCopilot is slightly higher, possibly due to successfully retrieving more required tables. We can also confirm the correlation between routing recall and EX by comparing these two baselines. CRUSHBM25 outperforms DTR in recall and EX on Spider, while DTR achieves higher EX on the other two datasets due to its superior recall. Therefore, the success of schema routing is essential for schema-agnostic NL2SQL.
Finding 5. DBCopilot outperforms retrieval baselines in schema-agnostic NL2SQL due to its superior schema routing.
Exploration of Prompt Strategies.  By generating multiple candidate schemas, DBCopilot can effectively increase the recall of the target schema. We explore three prompt strategies for leveraging multiple schemas. Interestingly, SQL generation with the top 1 schema performs best, even though the top 5 candidates have 22.44% more database recall on Spidersyn than the top 1. This may be due to confusion caused by similar candidate schemas, thus requiring more database/human information to disambiguate. By introducing human-in-the-loop (HITL) to select the desired schema from the top 5 candidates, DBCopilot comes close to the ideal situation of providing the gold database schema, but with lower inference costs. This highlights the ability of DBCopilot to exclude extraneous tables and its promise for schema-agnostic NL2SQL.
Finding 6. Using the best schema prompting strategy for SQL generation is optimal unless there is human interaction.
4.4. Ablation Study
Table 6. Performance loss in ablation studies. “BS”, “OD”, “CD”, “DB” stands for basic serialization, original NL2SQL training data, constrained decoding, and diverse beam search.
Spider Bird
Database Table Database Table
R@1 R@5 R@5 R@15 R@1 R@5 R@5 R@15
DBCopilot 85.01 96.42 91.63 97.51 88.92 97.13 85.83 94.59
   w/ BS -3.20 -2.90 -4.09 -5.75 -3.39 -1.63 -6.43 -5.99
   w/ OD -84.72 -82.40 -88.76 -79.55 -88.92 -82.27 -85.73 -88.81
   w/o CD -0.39 -0.19 -4.15 -2.51 +0.12 -0.45 -2.81 -1.93
   w/o DB +0.48 -3.00 +0.64 -2.88 -0.52 -4.89 -3.69 -8.87
We finally perform ablation studies on DBCopilot to validate the contribution of each module, as summarized in Table 6.
DFS Serialization.  DFS serialization is essential for schema routers to learn the underlying relations of databases and tables. To demonstrate this, we replace the DFS serialization with the unordered basic serialization of tables, similar to DSI for multiple documents. As shown in Table 6, this results in a 1% to 6% reduction in recall. DFS serialization effectively reduces the search space, whether by first generating a database and then generating its tables, or by continuing to generate based on existing tables. This graph-based serialization also makes constrained decoding possible.
Data Synthesis.  Generative retrieval is difficult to generalize to unseen documents during training. As we can see, routers trained on the original training instances ($N$, $S$) extracted from NL2SQL datasets are almost completely incapable of routing on test schemas, even when constrained decoding is used. Data synthesis by sampling schemas and generating pseudo-questions can effectively solve this challenge. Figure 6 shows how the performance varies with the amount of synthetic data. Database and table recalls increase rapidly at first as more data is synthesized, but level off after a certain amount. How to improve the quality of synthetic data and reduce the need for it may be a future direction.
Decoding Strategy.  Decoding optimizations (i.e., constrained decoding and group beam search) are beneficial for generative schema routing during inference. On the one hand, constrained decoding is useful for generating accurate table identifiers, since the drop from removing it is mainly in table recalls. On the other hand, group beam search helps to generate multiple diverse schemas because database recall@5 decreases significantly without it.
Finding 7. DFS serialization, data synthesis, and decoding optimizations contribute to DBCopilot in different ways.
5. Related Work
5.1. Natural Language to SQL
NL2SQL is an essential part of the natural language interface. Most research has approached this task along the lines of semantic parsing (Kamath and Das, 2019; Li et al., 2020). Early NL2SQL parsing systems adopted symbolic approaches that relied heavily on rule or grammar engineering, making them difficult to apply widely (Zelle and Mooney, 1996; Woods, 1973). In recent years, with the advances in deep learning, especially PLMs, end-to-end NL2SQL approaches that using neural generation models have become the mainstream (Li et al., 2023c; Wang et al., 2020; Gu et al., 2023a; Li et al., 2023a; Yu et al., 2021). Some efforts have been made to improve encoding and decoding, such as relation-aware encoding (Wang et al., 2020; Li et al., 2023a) and grammar-based decoding (Xu et al., 2017; Scholak et al., 2021). Some other efforts have focused on generating accurate SQL queries through schema linking (Li et al., 2023c; Wang et al., 2020) or skeleton generation (Gu et al., 2023a, b).
LLMs have emerged as a new paradigm for NL2SQL (Rajkumar et al., 2022; Liu et al., 2023a; Gao et al., 2023). Unlike previous work, the core of the LLM-based NL2SQL solution is prompt engineering, i.e., prompting LLM to generate correct and efficient SQL queries. Several efforts have been made to select appropriate in-context learning examples through question classification (Pourreza and Rafiei, 2023), skeleton retrieval (Guo et al., 2023), and in-domain example synthesis (Chang and Fosler-Lussier, 2023). Research has also been done on decoupling zero-shot NL2SQL into PLM-based sketch generation and LLM-based SQL completion (Gu et al., 2023b). These studies are orthogonal to our work and can be easily integrated to optimize DBCopilot SQL generation.
5.2. Table Discovery
Table discovery techniques are critical to data management, allowing users to explore and gain insights from data warehouses or lakes (Fan et al., 2023a). A common practice for table discovery is query-driven discovery (Zhang and Ives, 2020; Nargesian et al., 2019), where the query is typically in the form of keywords, tables, or NL questions. For keyword search, data discovery systems find relevant tables by searching for topic terms in the metadata (Cafarella et al., 2009; Pimplikar and Sarawagi, 2012). For table search, tables are retrieved to augment a given table as the query (Sarma et al., 2012). Joinable table search (Dong et al., 2021; Zhu et al., 2019) and unionable table search (Fan et al., 2023b; Khatiwada et al., 2023) aim to augment a query table with additional attributes and new tuples, respectively. For NL question search, open domain question answering retrieves independent tables to answer the question with table-based models (Chen et al., 2021; Herzig et al., 2021). Although dense retrieval methods based on contrastive learning have achieved advanced performance in the above tasks (Dong et al., 2021; Fan et al., 2023b; Herzig et al., 2021), they are difficult to consider the structural relationships between tables and retrieve them jointly. Alternatively, data navigation allows users to navigate interactively through the tables organized hierarchically (Nargesian et al., 2020) or relationally (Fernandez et al., 2018). Our work can be seen as automated data navigation, where we employ DSI to model the query-based navigation in an end-to-end manner.
Figure 6. Database recall@1 and table recall@5 w.r.t. the number of (#) synthetic training data.
6. Conclusion
In this paper, we present DBCopilot, a framework for effectively scaling natural language querying to massive databases. DBCopilot decouples schema-agnostic NL2SQL into schema routing and SQL generation through LLM-Copilot collaboration. We present an approach to effectively navigate the target database and tables for an NL question in a relation-aware, end-to-end manner. We also propose a training data synthesis paradigm to adapt routers over massive databases without manual intervention, and explore various prompt strategies to incorporate multiple candidate schemas. Extensive experiments verify the effectiveness of our solution. In summary, DBCopilot pushes the boundaries of NL2SQL, enabling researchers to better strategize their quest for data accessibility.
Appendix A Supplementary Experiments
A.1. Case Study
[…]
DPR $\langle\text{world},\{\text{countrylanguage}\}\rangle$
[…]
Figure 7. Successful case where DBCopilot routes the correct schema but baselines fail.
[…]
We provide successful cases to illustrate how DBCopilot improves schema routing through joint retrieval. As the example in Figure 7 shows, DBCopilot is the only approach that accurately routes the appropriate schema for the given natural language question. Sparse retrieval methods (BM25 and CRUSHBM25) rely heavily on lexical matching while ignoring the semantic matching between the question and the schemas. However, untrained dense retrieval methods (SXFMR and CRUSHSXFMR) also fail to capture the underlying semantics of tables and retrieve the wrong databases. Fine-tuning based approaches identify the correct database where DPR misses the country table linked to the countrylanguage table, but relation-aware DBCopilot does not.
[…]
A.2. Efficiency & Resource Consumption
We also measure the efficiency and resource consumption of each schema routing method. Table 7 shows the number of queries per second (QPS), the duration of build time, the amount of disk space utilized, and the consumption of system and GPU memory. LLM inference and complex post-processing are the main reasons for the slowness of CRUSH approaches. DBCopilot, based on an encoder-decoder architecture, consumes more disk space and GPU memory than dense retrieval approaches (SXFMR, DPR) based on an encoder-only architecture. Despite the slightly higher resource consumption for training and inference in generative retrieval, we believe and have observed that the growing popularity of generative AI is leading to an increasing number of optimizations being proposed.
Table 7. Efficiency & resource consumption of different methods. Build time includes the time for model training and index construction. We calculate queries per second (QPS) and memory consumption with a query batch size of 64.
Method QPS Build (s) Disk (MB) RAM (MB) GPU (MB)
BM25 1677.1 10 36 800 -
SXFMR 49.0 48 420 2,230 3742
CRUSHBM25 0.9 10 36 1104 -
CRUSHSXFMR 0.8 48 420 2502 3742
BM25 1677.1 6972 36 800 -
DPR 49.0 5161 420 2,230 3742
DBCopilot 43.0 13678 852 656 11807
A.3. Oracle Test with In-Context Learning
Table 8. Schema-Aware NL2SQL with In-Context Learning
Provided Schema Prompt Strategy Spider Bird Spidersyn
EX Cost EX Cost EX Cost
G. Column Basic 77.95 3.52 35.98 11.58 76.69 3.43
G. Table Basic 75.82 3.61 28.49 12.10 69.34 3.52
G. Database Basic 75.73 3.83 25.75 13.08 63.64 3.74
5 Database Basic 68.38 5.57 21.45 17.79 50.48 5.47
Table 8 shows the results of the oracle test by providing few-shot (5-shot) examples. As we can see, while this improves the accuracy of NL2SQL somewhat, it does not compensate for the performance degradation caused by extraneous schema elements. This makes schema linking still important in LLM-based NL2SQL. In the future, we will investigate how to better disambiguate and link schema elements in the context of schema-agnostic NL questions.

### v3

@@
-DBCopilot: Scaling Natural Language Querying to Massive Databases
+DBCopilot: Natural Language Querying over Massive Databases via Schema Routing
 Abstract.
-NL2SQL simplifies database interactions by enabling non-experts to convert natural language (NL) questions into Structured Query Language (SQL) queries.
-While recent advances in large language models (LLMs) have improved the zero/few-shot NL2SQL paradigm, existing methods face scalability challenges when dealing with massive databases.
-This paper introduces DBCopilot, a framework that addresses these challenges by employing a compact and flexible copilot model for routing over massive databases.
-Specifically, DBCopilot decouples schema-agnostic NL2SQL into domain-specific schema routing and generic SQL generation.
-This framework utilizes a lightweight differentiable search index to construct semantic mappings for massive database schemas and navigates natural language questions to their target databases and tables in a relation-aware, end-to-end manner.
-The routed schemas and questions are then fed into LLMs for effective SQL generation.
+The development of Natural Language Interfaces to Databases (NLIDBs) has been greatly advanced by the advent of large language models (LLMs), which provide an intuitive way to translate natural language (NL) questions into Structured Query Language (SQL) queries.
+While significant progress has been made in LLM-based NL2SQL, existing approaches face several challenges in real-world scenarios of natural language querying over massive databases.
+In this paper, we present DBCopilot, a framework that addresses these challenges by employing a compact and flexible copilot model for routing over massive databases.
+Specifically, DBCopilot decouples schema-agnostic NL2SQL into schema routing and SQL generation.
+This framework utilizes a single lightweight differentiable search index to construct semantic mappings for massive database schemata, and navigates natural language questions to their target databases and tables in a relation-aware joint retrieval manner.
+The routed schemata and questions are then fed into LLMs for effective SQL generation.
 Furthermore, DBCopilot introduces a reverse schema-to-question generation paradigm that can automatically learn and adapt the router over massive databases without manual intervention.
-Experimental results demonstrate that DBCopilot is a scalable and effective solution for schema-agnostic NL2SQL, providing a significant advance in handling large-scale schemas of real-world scenarios.
-Text-to-SQL, LLM, Generative Retrieval, Schema Routing
-copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXccs: Information systems Structured Query Languageccs: Information systems Top-k retrieval in databasesccs: Information systems Question answering
-Figure 1. The DBCopilot framework. DBCopilot decouples NL querying over massive databases into schema routing and SQL generation through LLM-Copilot collaboration.
+Experimental results verify that DBCopilot is a scalable and effective solution for schema-agnostic NL2SQL, providing a significant advance in handling natural language querying over massive databases for NLIDBs.
+copyright: rightsretaineddoi: isbn: 978-3-89318-099-8conference: 28th International Conference on Extending Database Technology (EDBT); 25th March-28th March, 2025; Barcelona, Spainjournalyear: 2025
 1. Introduction
-Natural language to SQL (NL2SQL) aims to convert natural language (NL) questions into Structured Query Language (SQL) queries, making it convenient for non-experts to query and analyze databases.
-This task plays a critical role in data analysis, business intelligence, and other database interaction applications, thus has attracted significant attention (Androutsopoulos et al., 1995; Affolter et al., 2019; Qin et al., 2022; Katsogiannis-Meimarakis and Koutrika, 2023; Fu et al., 2023; Gu et al., 2023a).
-The recent emergence of large language models (LLMs) brings a new zero/few-shot NL2SQL paradigm that directly transforms SQL queries via LLMs with schema-aware prompts, showing great promise (Rajkumar et al., 2022; Liu et al., 2023a; Pourreza and Rafiei, 2023; Gu et al., 2023b; Gao et al., 2023).
-However, existing work in this direction typically focuses only on querying a single database with a handful of tables, neglecting the real-world requirement to query over massive databases and data warehouses (Pride, 2023; Kothyari et al., 2023; Sen et al., 2020).
-In this case, data consumers need to invest significant effort to manually identify and select the appropriate database and relevant tables.
-Therefore, it becomes imperative to develop a mechanism that abstracts non-expert users from the intricacies of data representation and enables them to query over massive databases with schema-agnostic NL questions, i.e., without understanding of large-scale schemas.
+Natural Language Interfaces to Databases (NLIDBs) enable users to query and interact with databases by using natural language (NL) rather than complex query languages like Structured Query Language (SQL) (Androutsopoulos et al., 1995; Affolter et al., 2019).
+These systems are critical to data democratization by promoting data accessibility and are widely used in practical scenarios such as data analysis and business intelligence.
+The development of NLIDBs has been greatly advanced with the advent of large language models (LLMs), which provide an intuitive way to translate NL questions into SQL queries via schema-aware prompts in a new zero- or few-shot NL2SQL paradigm (Rajkumar et al., 2022; Liu et al., 2023; Pourreza and Rafiei, 2023; Gu et al., 2023b; Gao et al., 2024; Li et al., 2024c; Talaei et al., 2024; Li et al., 2024a).
+Figure 1.
+DBCopilot employs LLM-Copilot collaboration to scaling natural language querying over massive databases.
+Schema routing first navigates to the appropriate database and tables for NL questions via a copilot model.
+SQL generation then transforms SQL queries via LLMs with schema-aware prompts.
+However, existing work in this direction typically focuses only on querying a small database with a handful of tables, neglecting real-world scenarios of NLIDBs to query over massive databases (e.g., data lakes, data warehouses, and open data portals) with large-scale schemata (Pride, 2023; Kothyari et al., 2023; Sen et al., 2020; Stuart, 2015).
+As illustrated in Figure 1, data consumers in this case need to invest significant effort to manually identify and select the appropriate database and relevant tables.
+Therefore, it becomes imperative to develop a mechanism that abstracts non-experts from the complex data representation and enables them to query over massive databases with schema-agnostic NL questions, i.e., without understanding of large-scale schemata.
 The core of this process is to construct semantic mappings from diverse NL questions to massive database model elements, thus automatically identifying the target database and filtering for a minimal amount of relevant tables.
-The above scenarios pose significant challenges to current LLM-based NL2SQL methods.
-Due to token limits and inference costs, it is infeasible to feed hundreds/thousands of schemas into LLMs for each question.
-Furthermore, current LLMs still struggle to effectively leverage information present in long contexts (Liu et al., 2023b; Xu et al., 2023).
-To address this scalability problem, existing solutions can be categorized into retrieval-based and fine-tuning based methods.
-Given an NL question, retrieval-based methods identify the corresponding table schemas using information retrieval techniques (Wikipedia contributors, 2023a; Karpukhin et al., 2020).
-Unfortunately, by relying on the co-occurrence of words and representations between question and schema, retrieval-based methods suffer from vocabulary mismatch and the diversity of NL expressions.
-In addition, these ad hoc retrieval methods typically treat documents as independent objects, neglecting the relationships between databases and tables.
-Fine-tuning based methods inject schema knowledge into the parameters of LLMs by fine-tuning with schema descriptions or (question, schema) pairs.
-Unfortunately, fine-tuning LLMs with billions of parameters to keep up with evolving database schemas is resource-intensive and sometimes impractical.
-Besides, it remains challenging to effectively inject new knowledge into LLMs through fine-tuning (Ovadia et al., 2023).
-As a result, how to extend the scalability of LLM-based NL2SQL methods and apply them to massive databases is a major challenge in this field.
+Limitations of state-of-the-art.  The above scenarios pose significant challenges to current LLM-based NL2SQL, schema linking, and ad-hoc retrieval approaches, including:
+  1. C1)
+    Schema Scalability: It is infeasible to feed hundreds or thousands of schemata into LLMs for each question due to token limits and inference costs.
+Moreover, current LLMs struggle to effectively leverage information present in long contexts (Liu et al., 2024; Levy et al., 2024).
+Therefore, numerous studies have employed schema linking as a prerequisite step for LLM-based NL2SQL (Pourreza and Rafiei, 2023; Wang et al., 2025; Li et al., 2024c).
+However, they follow the typical practice of simultaneously feeding the entire schema and NL question into a model (Dong et al., 2019; Lei et al., 2020; Li et al., 2023c), which hinders their scalability when dealing with large-scale schemata.
+  2. C2)
+    Schema Complexity: Real-world database schemata are characterized by inherent intricacy and heterogeneity in the relationships between tables, as shown in Figure 1.
+These relationships are necessary to identify the correct schema because SQL queries for NL questions often span multiple related tables.
+However, ad-hoc retrieval methods such as BM25 (Wikipedia contributors, 2023b) and DPR (Karpukhin et al., 2020) for schema linking or retrieval-augmented generation (RAG) (Zhao et al., 2024; Gao et al., 2023) merely retrieve elements individually.
+This approach is suboptimal because the relationships can only be considered through post-processing or re-ranking (Kothyari et al., 2023; Chen et al., 2024) rather than in an end-to-end manner.
+  3. C3)
+    Semantic Mismatch: Unlike general natural language text, database schemata often employ domain-specific or unconventional terminology and naming conventions for tables and columns (Floratou et al., 2024).
+This discrepancy entails a semantic mismatch between NL questions and schemata.
+As a result, retrieval-based methods that rely on the co-occurrence of words or sentence representations suffer from the vocabulary mismatch and the diversity of schemata (Talaei et al., 2024; Li et al., 2024c).
+While some efforts attempt to rewrite NL questions into possible schema elements via LLMs for format alignment (Kothyari et al., 2023; Chen et al., 2024), they do not adequately address the underlying semantic mismatch.
+Approach.
 In this paper, we present DBCopilot, a framework for effectively scaling natural language querying to massive databases.
-The main idea behind DBCopilot is to introduce a compact and flexible copilot model to serve as a router across massive databases, and to leverage the efficient NL2SQL capability of LLMs to generate final SQL queries.
-To this end, as shown in Figure 1, DBCopilot decouples schema-agnostic NL2SQL into domain-specific schema routing and generic SQL generation.
-Specifically, DBCopilot utilizes a lightweight differentiable search index (DSI) (Tay et al., 2022) to navigate an NL language question to its corresponding database and tables.
-Then the routed schemas and questions are fed into LLMs to generate SQL queries by leveraging their strong language understanding and SQL generation capabilities.
-In this way, DBCopilot can easily integrate with advanced LLM-based NL2SQL solutions and rapidly and cost-effectively scale to massive databases.
-For an effective implementation of DBCopilot, three major challenges need to be addressed.
-(1) Schema routing is not a task of retrieving independent tables because complex queries require information across multiple tables, and there are inherent relationships between databases and tables.
-To resolve this challenge, we model schema routing as a sequence-to-sequence (Seq2Seq) DSI.
-By proposing schema graph construction, DFS serialization, and constrained decoding, we can jointly retrieve target databases and tables in an end-to-end manner, taking into account structural relationships.
-(2) The generative retrieval paradigm makes it infeasible to perform semantic mapping of unseen schemas because they are not indexed into model parameters.
-In practice, labeling NL questions for massive database schemas is highly expensive.
-To this end, we propose a reverse schema-to-question generation paradigm to automatically synthesize training data and efficiently train the schema router.
-(3) How to incorporate routed schemas with LLMs to perform schema-aware SQL generation is critical to the quality of NL2SQL.
-We explore various prompt strategies to select and incorporate multiple candidate schemas for DBCopilot.
-We use three standard NL2SQL datasets and two additional robustness datasets to evaluate DBCopilot.
-Experimental results verify the effectiveness of DBCopilot, with its copilot model outperforming retrieval-based baselines in schema routing by up to 19.88% in recall, and its schema-agnostic NL2SQL showing an improvement in execution accuracy of more than 7.35%.
-In summary, we demonstrate that DBCopilot is a scalable and effective solution for schema-agnostic NL2SQL, presenting a significant step forward in handling large-scale schemas of real-world scenarios.
-Contributions.  Our contributions can be summarized as follows.
-  • We highlight the limitations of existing NL2SQL methods in scaling from limited tables to massive databases.
-  • We present DBCopilot, a framework that decouples schema-agnostic NL2SQL into schema routing and SQL generation via an LLM-copilot collaboration architecture.
-  • We propose a relation-aware, end-to-end joint retrieval approach for schema routing, and also propose a reverse generation paradigm for automatic training data synthesis.
-  • We conduct thorough experiments to verify the effectiveness of our solution in scaling NL2SQL to massive databases.
+The main idea behind DBCopilot is to introduce a compact and flexible copilot model for routing over massive databases, and to leverage a powerful LLM for generating SQL queries.
+By decoupling schema-agnostic NL2SQL into schema routing and SQL generation, DBCopilot avoids frequently adapting LLMs to domain-specific data while maintaining their generic SQL generation capability.
+As shown in Figure 1, DBCopilot utilizes a lightweight Differentiable Search Index (DSI) (Tay et al., 2022) to navigate an NL question to its corresponding database and tables.
+Then the routed schemata and questions are fed into LLMs to generate SQL queries by leveraging their strong capabilities.
+In this way, DBCopilot can be easily integrated with advanced LLM-based NL2SQL solutions and scale rapidly and cost-effectively to massive databases for NLIDBs.
+To effectively implement DBCopilot, we address the above challenges in the following way.
+For C1) schema scalability, we leverage generative retrieval (a.k.a.
+DSI) to avoid the need to simultaneously feed whole schemata and the NL question into a neural model.
+By injecting schema knowledge into the parameters of the schema router, the corresponding SQL query schema can be generated by simply inputting an NL question, thereby extending scalability.
+For C2) schema complexity, we propose a relation-aware joint retrieval approach to identify the target databases and tables in an end-to-end manner.
+We first construct a schema graph to represent the relationships between databases and tables.
+Based on this graph, we design DFS serialization and constrained decoding algorithms to facilitate relation-aware Seq2Seq retrieval.
+For C3) semantic mismatch, we introduce a reverse schema-to-question generation paradigm to automatically synthesize mass training data and efficiently learn the semantic mapping for the schema router.
+This approach effectively mitigates the labeling cost for massive database schemata and the problem that generative retrieval cannot generalize to unseen schemata during training.
+Finally, we explore various prompt strategies to select and incorporate multiple candidate schemata for LLM-based SQL generation.
+Evaluation.
+We adapt three standard NL2SQL datasets and two robustness datasets to evaluate DBCopilot.
+Experimental results verify the effectiveness of DBCopilot, with its copilot model outperforming retrieval-based baselines in schema routing by up to 19.88% in recall, and its schema-agnostic NL2SQL showing an improvement in execution accuracy of 4.43%~11.22%.
+In summary, we demonstrate that DBCopilot is a scalable and effective solution for schema-agnostic NL2SQL, presenting a significant step forward in handling large-scale schemata of real-world scenarios.
+Contributions.  Our contributions are summarized as follows. [^1] We have open-sourced our code and data at: github.com/tshu-w/DBCopilot.
+  • We approach the practical and challenging problem of NLIDBs for querying over massive databases, while pointing out the limitations of existing approaches.
+  • We present DBCopilot, a framework that addresses this problem by decoupling schema-agnostic NL2SQL into schema routing and SQL generation via LLM-copilot collaboration.
+  • We propose a relation-aware joint retrieval approach for end-to-end schema routing and a reverse generation paradigm for learning semantic mappings via training data synthesis.
+  • We conduct thorough experiments to verify the effectiveness of DBCopilot in scaling NL2SQL to massive databases.
 2. Preliminaries
-2.1. Problem Definition
+2.1. Problem Formulation
 Table 1. Notation list used in this paper.
-Symbol Description
+[…]
+notes/style=0 {NiceTabularX} 1.13cX Symbol Description
 $N$ Natural language question
@@
 $T$ Set of tables involved in query $Q$, with their properties
-$S$ SQL query schema $\langle D,T\rangle$, where $T$ is the subset of tables belonging to database $D$
+$S$ SQL query schema $\langle D,T\rangle$, where $T$ is the subset of tables in database $D$
 $\mathcal{T}$ Union of sets of tables from each database in $\mathcal{D}$
-$\mathcal{S}$ Schema space $\mathcal{D}\times\mathcal{P}(\mathcal{T})$, where $\mathcal{P}(\mathcal{T})$ is the power set of $\mathcal{T}$
+[…]
 $\mathcal{G}$ Schema graph $\langle\mathcal{V},\mathcal{E}\rangle$, where $\mathcal{V},\mathcal{E}$ are graph nodes and edges
-As mentioned above, we formulate the problem of natural language querying over massive databases and large numbers of tables with unknown SQL query schema (i.e., target database and tables) as schema-agnostic NL2SQL.
-Formally, as per the notation list detailed in Table 1, the problem of schema-agnostic NL2SQL is defined in Definition 2.1, and its two subtasks schema routing and SQL generation are defined in Definition 2.2 and Definition 2.3, respectively.
-Obviously, traditional NL2SQL with an NL question and its limited database schema as input is similar to the subtask SQL generation.
+As mentioned above, we abstract and formulate the problem of natural language querying over massive databases and large numbers of tables with unspecified SQL query schema (i.e., the target database and tables) as schema-agnostic NL2SQL.
+Formally, as per the notation list detailed in § 2.1, the problem of schema-agnostic NL2SQL and its two subtasks schema routing and SQL generation are described in Definition 2.1, 2.2, and 2.3, respectively.
 Definition 0 (Schema-Agnostic NL2SQL).
-Given a natural language question $N$, schema-agnostic NL2SQL generates an equivalent SQL query $Q$ that can be executed against the appropriate database $D$ drawn from a collection of databases $\mathcal{D}$, without being provided with the intended SQL query schema $S=\langle D,T\rangle$.
+Given a natural language question $N$, schema-agnostic NL2SQL generates an equivalent SQL query $Q$ that can be executed against the appropriate database $D$ drawn from a set of databases $\mathcal{D}$, without being provided with the intended SQL query schema $S=\langle D,T\rangle$.
 Definition 0 (Schema Routing).
-Schema routing is the process of identifying the SQL query schema $S=\langle D,T\rangle$ of a natural language question $N$ from the entire space of all possible schemas $\mathcal{S}=\mathcal{D}\times\mathcal{P}(\mathcal{T})$ within the specified set of databases $\mathcal{D}$ and tables $\mathcal{T}$.
+Schema routing identifies the SQL query schema $S=\langle D,T\rangle$ of a natural language question $N$ from the space of all possible schemata $\mathcal{S}=\mathcal{D}\times\mathcal{P}(\mathcal{T})$ within the specified set of databases $\mathcal{D}$ and tables $\mathcal{T}$.
 Definition 0 (SQL Generation).
@@
 Schema Routing.
-Schema linking (Katsogiannis-Meimarakis and Koutrika, 2023) is an optional substep of traditional NL2SQL methods.
-Given an NL question and a database schema, schema linking aligns NL terminologies, called “mentions”, with their corresponding schema elements (such as tables, columns, or database values).
+Schema linking is an optional substep of NL2SQL methods (Katsogiannis-Meimarakis and Koutrika, 2023).
+Given an NL question and a database schema, schema linking aligns natural language terminologies, called “mentions”, with their corresponding schema elements (such as tables, columns, and database values).
 In this way, schema linking servers as a bridge between NL questions and database elements.
-Schema linking and schema routing can be distinguished in two aspects.
-First, schema linking is a substep of SQL generation for accurate generation, while schema routing is a prerequisite of SQL generation to identify the unknown schema of a question.
-Second, the input for schema linking is the NL question along with the schema, whereas, the input for schema routing is the NL question and the output is its indexed or memorized schema.
+Schema routing is an extension of the schema linking concept adapted for querying over massive databases.
+Specifically, schema routing differs from existing schema linking methods in the following aspects.
+First, schema linking methods (Dong et al., 2019; Lei et al., 2020; Li et al., 2023c) usually take the entire single-database schema and an NL question as input, while schema routing, driven by scalability requirements, should take the NL question as input with massive database schemata indexed.
+Second, schema routing care about coarse-grained databases and tables, while leaving the fine-grained columns and values for further SQL generation.
 2.2. Generative Retrieval
-Traditional information retrieval (IR) techniques typically follow the “index-retrieve” paradigm.
-However, such a pipelined framework, which cannot be optimized together in an end-to-end manner, limits the final retrieval quality.
+Traditional information retrieval (IR) techniques follow the “index-retrieve” paradigm.
+Such a pipeline, which cannot be optimized together in an end-to-end manner, limits the final retrieval quality.
 To address this concern, generative retrieval (Tay et al., 2022; Wang et al., 2022) (a.k.a.
-DSI) is an emerging paradigm that trains a Seq2Seq pre-trained language model (PLM) to map queries directly to relevant document identifiers.
-The main idea behind it is to completely parameterize traditional “index-retrieve” pipelines within a single neural model, by modeling indexing and retrieval operations into the model parameters.
-Critical ongoing research in this paradigm focuses on defining better document identifiers (Sun et al., 2023; Zhang et al., 2023).
-To the best of our knowledge, this paper is the first to adapt this paradigm to jointly retrieve multiple elements with structural relationships.
+DSI) is an emerging paradigm that leverages a Seq2Seq pre-trained language model (PLM) to directly generate relevant document IDs (e.g., “doc456”) in response to input queries, rather than vector representations for computing similarities.
+Technically, DSI works by Seq2Seq learning on (query/doc_chunk, doc_id) pairs to model semantic mappings, completely parameterizing traditional “index-retrieve” pipelines within a single neural model.
+Therefore, unlike conventional machine learning tasks that aim to generalize to unseen data, the training phase of DSI shifts to memorize all retrieval targets – including those used in evaluation – similar to the indexing phase of traditional retrieval.
+Critical ongoing research focuses on defining better document IDs (Sun et al., 2023; Zhang et al., 2023).
+However, these studies mainly focus on retrieving elements individually without considering their relationships (e.g., joinable relations between tables, triple relations in knowledge graphs, or semantic relations between documents).
+To the best of our knowledge, this paper is the first to adapt this paradigm to jointly retrieve elements with relationships, opening up new opportunities for applying it in graph-based retrieval (Peng et al., 2024).
+2.3. LLM-based Natural Language to SQL
+LLMs, developed by scaling up PLMs in terms of both model and data size, have emerged as a milestone in artificial intelligence (Zhao et al., 2023).
+Despite their similar architecture and pre-training tasks, LLMs exhibit unexpected emergent abilities (e.g., in-context reasoning, instruction following) compared to small PLMs (Wei et al., 2022).
+These emergent abilities change the paradigm of model use from “fine-tuning” to “prompt engineering” that does not change the model parameters.
+This shift has brought new opportunities for all kinds of natural language processing tasks, including NL2SQL.
+Coupled with well-designed schema-aware prompts and few-shot NL2SQL examples, LLMs can generate SQL queries directly from NL questions, delivering state-of-the-art performance (Gao et al., 2024; Gu et al., 2023b; Li et al., 2024c).
+In addition, pipeline-based approaches decompose NL2SQL into substeps including schema linking, query classification, query decomposition, SQL refinement, further extending the power of LLM-based NL2SQL (Pourreza and Rafiei, 2023; Li et al., 2024b).
+Nevertheless, LLM-based NL2SQL still faces significant challenges in real-world scenarios, such as schema scalability, complex or ambiguous queries, SQL efficiency, and so on (Floratou et al., 2024; Li et al., 2023b; Kothyari et al., 2023).
 3. Methodology
-This section introduces DBCopilot, a schema-agnostic NL2SQL framework for effectively scaling NL querying to massive databases.
+This section introduces DBCopilot, a schema-agnostic NL2SQL framework for effectively scaling natural language querying to massive databases.
+Figure 2.
+Training process of schema router.
+We first construct a schema graph to represent the relationships of all databases and their tables.
+Based on this graph, we sample SQL query schemata and generate pseduo-questions using a reverse-trained schema questioning model that synthesizes training data.
+Finally, we apply DFS serialization and train the schema router.
 3.1. Overview
-As shown in Figure 1, DBCopilot decouples natural language querying over massive databases into schema routing and SQL generation through LLM-Copilot collaboration.
+As shown in Figure 1, DBCopilot decouples natural language querying over massive databases into schema routing and SQL generation.
 Specifically, given an NL question $N$, DBCopilot utilizes a copilot model as a schema router to identify the SQL query schema $S=\langle D,T\rangle$ associated with that question.
@@
 Example 3.1 provides an illustration of this process.
+Consequently, this LLM-Copilot collaboration architecture enables easy integration with advanced LLM-based NL2SQL solutions and cost-effective scaling to massive databases for NLIDBs.
 Example 0.
-Given a question “Which singers held concerts in 2022?”, DBCopilot converts it into an equivalent SQL query through the following two phases.
-1. Schema Routing identifies the relevant database and tables that contain the information required to answer the question:
+Given a natural language question “Which singers held concerts in 2022?”, DBCopilot converts it into an equivalent SQL query through the following two phases:
+  1. 1.
+    Schema Routing identifies the relevant database and tables containing necessary information to answer the question.
 […]
-2. SQL Generation prompts the LLM with the identified schema and question, and generates the SQL query:
-SELECT s.name FROM singer_in_concert AS sic JOIN singer AS s
-ON sic.singer_id = s.singer_id JOIN concert AS c
-ON sic.concert_id = c.concert_id WHERE c.year = 2022
-Figure 2. The training process of the schema router.
-In practice, the schema space $\mathcal{S}=\mathcal{D}\times\mathcal{P}(\mathcal{T})$ is huge because it is the cartesian product of the database collections $\mathcal{D}$ and the power set of all tables $\mathcal{T}$.
-To narrow down the search space, we propose a relation-aware, end-to-end joint retrieval approach for schema routing.
-The training process is illustrated in Figure 2.
-Specifically, we first construct a schema graph to represent the relationships between databases and tables (§ 3.2).
+  2. 2.
+    SQL Generation prompts LLMs with the identified schema and NL question to generate the SQL query.
+    SELECT s.NAME
+    FROM concert_singer.singer_in_concert AS sic
+    JOIN concert_singer.singer AS s
+    ON sic.singer_id = s.singer_id
+    JOIN concert_singer.concert AS c
+    ON sic.concert_id = c.concert_id
+    WHERE c.year = 2022
+To better scale schema routing to massive databases by modeling relationships between schema elements and learning semantic mappings from natural language to schemata, we propose a relation-aware joint retrieval approach for end-to-end schema routing.
+Specifically, we first construct a schema graph to represent the underlying relationships between databases and tables (§ 3.2).
 Based on this graph, we design a schema serialization algorithm to serialize a SQL query schema into a token sequence so that it can be generated by Seq2Seq DSI (§ 3.3).
-To address the challenge that DSI cannot generalize to unseen schemas, we further propose a reverse generation paradigm for training data synthesis (§ 3.4).
-These measures allow us to effectively train the schema router using serialized pseudo-instances ¡$N$, $S$¿ and apply inference with graph-based constrained decoding (§ 3.5).
-Finally, we explore various prompt strategies to select and incorporate multiple candidate schemas with LLMs (§ 3.6).
+To bridge the semantic gap between natural language and domain-specific schemata, we further propose a reverse generation paradigm for training data synthesis (§ 3.4).
+These measures allow us to effectively train the schema router with serialized pseudo-instances <$N$, $S$> and apply inference with graph-based constrained decoding (§ 3.5).
+Figure 2 illustrates the training process.
+Finally, we explore various prompt strategies to select and incorporate multiple candidate schemata for LLM-based SQL generation (§ 3.6).
 3.2. Schema Graph Construction
-Unlike NL expressions, a valid SQL query must follow the structure of the database.
-Therefore, it is necessary to inform the schema router about the inherent relationships/constraints between schema elements.
-In general, the schema $S=\langle D,T\rangle$ involved in a valid SQL query must satisfy the following conditions.
-First, the set of tables $T$ must be the subset of tables belonging to database $D$, since a SQL query obviously cannot reference tables that do not exist in the target database.
-Second, the tables in $T$ should cross-reference at least one other table with a common column (e.g., foreign key), so that the SQL query can link them together via join or subquery clauses.
-In addition to explicit Primary-Foreign relations, two tables can also be linked by implicit Foreign-Foreign relations, i.e., they reference the same unique key column of another table.
+Unlike natural language expressions, a valid SQL query must follow the logical data model (Wikipedia contributors, 2023a) of the database.
+Therefore, it is necessary to inform the schema router about the inherent relationships and constraints between schema elements.
+In general, the SQL query schema $S=\langle D,T\rangle$ involved in a valid SQL query should satisfy the following conditions.
+First, the set of tables $T$ should be the subset of tables belonging to database $D$, since a SQL query should not reference tables that do not exist in the target database.
+Second, the tables in $T$ should cross-reference at least one other table with a common column (e.g., foreign key), so that the SQL query can connect them together via join or subquery clauses.
+Example 3.2 and 3.3 show two examples.
+If two tables reference different columns of the junction table, then this junction table is required in a SQL query for these two tables.
+However, if two tables reference the same column of another table, they can be linked implicitly.
 Example 0.
-Given a question “Which rivers run through the state with the largest city in the US?”, the corresponding SQL query “SELECT river_name FROM river WHERE traverse IN ( SELECT state_name FROM city WHERE population = ( SELECT MAX ( population ) FROM city ) )” uses only the city and river tables, both of which refer to the primary key of the state table.
-To this end, we construct a directed graph $\mathcal{G}=\langle\mathcal{V},\mathcal{E}\rangle$ to represent the underlying schema structure of all databases $\mathcal{D}$ and their tables $\mathcal{T}$.
-Concretely, the schema graph consists of three types of nodes $\mathcal{V}=\{\nu_{s}\}\cup\mathcal{D}\cup\mathcal{T}$, where $\nu_{s}$ is a special node denoting the set of all databases $\mathcal{D}$.
-Each edge $e\in\mathcal{E}$ in this graph represents a relation between two nodes, which can be divided into:
-  • Inclusion relations: indicate whether a database is part of a database collection and whether a table belongs to a database.
-  • Table relations: include explicit Primary-Foreign and implicit Foreign-Foreign relations between table nodes.
+Given a natural language question “What are the names of the singers who performed in a concert in 2014?”, the corresponding SQL query is:
+SELECT s.name
+FROM singer_in_concert AS sc
+JOIN singer AS s ON sc.singer_id = s.singer_id
+JOIN concert AS c ON sc.concert_id = c.concert_id
+WHERE c.year = 2014
+This query involves the singer, concert, and singer_in_concert tables, where singer_in_concert is the join table with two different columns referring to singer and concert respectively.
+Example 0.
+Given a natural language question “Which rivers run through the state with the largest city in the US?”, the corresponding SQL query is:
+SELECT river_name FROM river
+WHERE traverse IN (
+SELECT state_name FROM city
+WHERE population = (
+SELECT MAX(population) FROM city
+This query involves only the city and river tables, both of which refer to the primary key of the state table.
+To this end, we construct a directed graph $\mathcal{G}=\langle\mathcal{V},\mathcal{E}\rangle$ to represent the underlying schemata of all databases $\mathcal{D}$ and their tables $\mathcal{T}$.
+Concretely, the schema graph consists of three types of nodes $\mathcal{V}=\{\nu_{s}\}\cup\mathcal{D}\cup\mathcal{T}$ and many types of edges =$\mathcal{E}=\{e_{1},\ldots,e_{|\mathcal{E}|}\}$, where $\nu_{s}$ is a special node denoting the set of all databases $\mathcal{D}$.
+Each edge $e\in\mathcal{E}$ in this graph represents a relation between two nodes that can be divided into two categories:
+  • Inclusion relations: Inclusion relations indicate whether a database is part of the database collection and whether a table belongs to a database.
+  • Table relations: Table relations include explicit Primary-Foreign, implicit Foreign-Foreign, and Joinable relations between table nodes.
+Joinable (Sarma et al., 2012) indicates that two tables share some values in certain attributes, allowing them to be combined together in SQL queries.
 1
-Input: A dictionary of all database schemas $databases$.
+Input: A dictionary of all database schemata $databases$.
 Output: The heterogeneous directed schema graph $\mathcal{G}$.
@@
 7             $\mathcal{G}$.addEdge($database$, $table$)
-8             for $column$ in $table$ :
-9                   if $column$ contains a foreign key :
-10                         $ltbl$, $lcol$leftarrow $column.{\mkern-2.0muf\mkern-2.0mu}oreign\_key$
-11                         Add $ltbl$ and $table$ into set $links[ltbl.lcol]$
+             // Joinable covers the cases of Primary-Foreign and Foreign-Foreign.
+8             for $relatedTable$ in getJoinableTables($table$) :
+9                   $\mathcal{G}$.addEdge($table$, $relatedTable$))
+10                   $\mathcal{G}$.addEdge($relatedTable$, $table$))
+11
 12
@@
 15
-16
-17
-18      for $tables$ in $links$ :
-19             for $tbl1$, $tbl2$ in $tables\times tables$ :
-20                   if $tbl1$ is not equal to $tbl2$ :
-21                         $\mathcal{G}$.addEdge($tbl1$, $tbl2$)
-22
-23
-24
-25
-26
-27
-28
-29
-30return $\mathcal{G}$
+16return $\mathcal{G}$
 Algorithm 1 Schema Graph Construction.
-Algorithm 1 shows the procedure for constructing this three-tiered hierarchical schema graph.
-By capturing the structure of database schemas in graph $\mathcal{G}$, any valid SQL query schema $S$ is a trail (Wikipedia contributors, 2023b) on this graph.
+Algorithm 1 shows the procedure for constructing this three-tiered hierarchical graph.
+By capturing the logical structure of database schemata in graph $\mathcal{G}$, any valid single-database SQL query schema $S$ is a trail (Wikipedia contributors, 2023c) on this graph.
 Each schema trail starts with node $\nu_{s}$, followed by a database node and some of its interconnected table nodes.
-The schema graph can efficiently narrow down the schema space $\mathcal{S}$ into trials on this graph starting from node $\nu_{s}$.
-As a result, the schema graph is essential for effectively representing and sampling SQL query schemas, serving as the foundation of DBCopilot.
+In this way, the schema graph facilitates relation-based serialization, valid SQL query schema sampling, and graph-based constrained decoding introduced below, serving as the foundation of DBCopilot.
 3.3. Schema Serialization
-Since we model schema routing as a Seq2Seq generative retrieval process, we need to design a schema serialization algorithm that converts a SQL query schema into a sequence of tokens.
+In this paper, we model schema routing as a generative retrieval process from an NL question $N$ to a SQL query schema $\langle D,T\rangle$.
+Since a Seq2Seq LM takes a sequence of tokens as input and outputs a sequence of tokens, we need to design a schema serialization algorithm that converts the partially ordered SQL query schema into a sequence of tokens.
 Basic Serialization.
-A trivial solution is to randomly order the set of tables as ad hoc retrieval (Chen et al., 2023), and identify the database based on the votes of the generated tables.
-However, this method loses sight of the inclusion and table relations between schema elements.
-The former results in a huge search space due to the lack of database constraints during table generation.
-The latter can result in missing tables during schema routing, especially junction tables.
+A trivial solution is to randomly order the set of tables as DSI for ad-hoc retrieval (Chen et al., 2023) and identify the target database based on the votes of the generated tables after inference.
+However, this serialization loses sight of the inclusion and table relations between schema elements.
+Due to the position bias and autoregressive nature of Seq2Seq LMs, the order of schema elements is important for training routers to better model their relationships (Liu et al., 2024; Berglund et al., 2024).
+Therefore, we need a carefully designed serialization strategy that preserves relationships within SQL query schema.
 Depth-First Search Serialization.
@@
 $\operatorname{Serialize}_{\pi}(S)\coloneqq\operatorname{Concat}\left(\left[\operatorname{DFS_{\pi}}(S,\mathcal{G})[i]\right]_{i=1}^{\left|\operatorname{DFS_{\pi}}(S,\mathcal{G})\right|}\right)$
-where $\pi$ is a node iteration order and $\operatorname{DFS_{\pi}}(S,\mathcal{G})$ is the DFS traversal sequence of schema $S$ on schema graph $\mathcal{G}$ when iterating successor nodes under order $\pi$.
-We introduce the iteration order parameter $\pi$ to determine the DFS sequence because a SQL query schema is not a completely ordered set and there may be multiple DFS sequences for a schema.
-We randomly select one DFS sequence at a time as the model target during training and assign different probabilities to different sequences originating from the same schema during inference.
+where $\pi$ is the node iteration order and $\operatorname{DFS_{\pi}}(S,\mathcal{G})$ is the DFS traversal sequence of schema $S$ on schema graph $\mathcal{G}$ when iterating successor nodes under order $\pi$ and $[i]$ indicates the $i$-th node of the traversal sequence.
+We introduce the iteration order parameter $\pi$ to determine the DFS sequence because a SQL query schema is not a completely ordered set and there may be multiple DFS sequences for a SQL query schema.
+We randomly select one DFS sequence at a time as the model target during training, and the schema router assigns different probabilities to different sequences originating from the same schema during inference.
 1
-Input: An ordered pair of a SQL query schema $S=\langle D,T\rangle$, the constructed schema graph $\mathcal{G}$, and a node iteration order $\pi$.
+Input: An ordered pair of a SQL query schema $S=\langle D,T\rangle$, the schema graph $\mathcal{G}$, and a node iteration order $\pi$.
 Output: The serialization of schema $S$.
@@
 10      $successors$leftarrow [][\;][ ]
+       // Iterates over successor nodes according to node iteration order $\pi$
 11       for $successor$ in $\mathcal{G}$[$node$] :
-             /* Iterates over successor nodes according to node iteration order $\pi$/
 12             if $successor$ in $nodes$ and not in $visited$ :
@@
 3.4. Training Data Synthesis
-Based on the schema graph and the DFS serialization, we can learn the schema router on question-schema pairs ($N$, $S$) as described in § 3.5.
-However, the generative retrieval paradigm prevents the schema router from generalizing to unseen schemas because they are not injected into the model parameters.
-To address this challenge, we propose a labor-free training data synthesis method that automatically generates pseudo-instances via a reverse schema-to-question generation paradigm, as shown in Figure 2.
-Specifically, we sample a sufficient number of valid schemas from the schema graph and generate a pseudo-question for each sampled schema.
-These schemas are sampled by performing finite-length random walks over the schema graph, and the database and tables traversed by a walk form a sampled schema.
+Based on the schema graph and DFS serialization, we can train the schema router on question-schema pairs ($N$, $S$) as described in § 3.5.
+However, it is infeasible to manually label training data for massive database schemata.
+Furthermore, the generative retrieval paradigm prevents the schema router from generalizing to unseen schemata because they are not injected into model parameters.
+To address these challenges, we introduce a labor-free training data synthesis method that automatically generates pseudo-instances via a reverse schema-to-question generation paradigm, as illustrated in Figure 2.
+Specifically, we sample a large number of valid SQL query schemata from the schema graph and generate a pseudo-question for each sampled schema.
+These schemata are sampled by performing finite-length random walks starting at node $\nu_{s}$ over the schema graph, and the database and tables traversed by a walk form a sampled schema.
 To generate an NL question for each sampled schema, we train a schema questioning model $\mathcal{M}_{\text{q}}$ in reverse based on the training sets of NL2SQL datasets.
+The reason for training a specific questioning model rather than using LLMs directly is to save the cost of synthesizing large amounts of training data.
 We use a SQL parser to extract the metadata (tables and columns) of SQL queries from the training pairs ($N$, $Q$).
-Unlike the schema router, which takes a condensed SQL query schema as output, the schema questioner accepts more detailed schema information (such as table columns) as input to generate rich and detailed questions.
+Unlike the schema router, which takes a condensed SQL query schema as output, the schema questioner accepts more detailed schema (e.g., table columns and comments) as input to generate rich and detailed questions.
 Given an illustration of the model input and output in Figure 3, the training loss of the schema questioning model $\mathcal{M}_{q}$ with parameters $\theta_{\text{q}}$ on $m$ extracted schema-question pairs ($S$, $N$) is:
 $\mathcal{L}_{q}=-\sum_{i=1}^{m}\log P\left(N_{i}\mid S_{i};\theta_{\text{q}}\right)$
+In practice, it can also generate high-quality seed data from schemata through LLMs and further boost them in this way.
 Pseudo-Question Generation Input: Ask a question that needs to be answered by combining the contents of all the tables, based on the word1 database schema provided below.
-- countrylanguage(language, countrycode, official) - country(code, lifeexpectancy) Output: What is the official language of the country with the highest life expectancy?
+- countrylanguage(language, countrycode, official) [Optional schema comments, if available.] - country(code, lifeexpectancy) Output: What is the official language of the country with the highest life expectancy?
 Figure 3. An illustration of pseudo-question generation.
-3.5. Schema Router Training & Inference
-Through the above measures, we can efficiently train the schema router to parameterize the semantic mapping of massive database schemas and effectively route NL questions to their target schemas via graph-based constrained decoding.
-Formally, the Seq2Seq schema router $\mathcal{M}_{r}$ takes a question sequence $N=\left[t_{1},t_{2},\ldots,t_{|N|}\right]$ as input and generates the serialized schema sequence $\bar{S}=\left[s_{1},s_{2},\ldots,s_{|\mathbf{s}|}\right]$.
+3.5. Schema Router Training & Constrained Inference
+Through the above measures, we can efficiently train the schema router to parameterize the semantic mapping of massive database schemata and effectively route NL questions to their target schemata via graph-based constrained decoding.
+Formally, the Seq2Seq schema router $\mathcal{M}_{r}$ takes a token sequence of question $N$ as input and generates the serialized schema sequence $\bar{S}$.
 Suppose the synthetic data is $\{(N_{i},S_{i})\mid 1\leq i\leq n\}$, the training loss of schema router $\mathcal{M}_{r}$ with parameters $\theta_{r}$ is:
 $\mathcal{L}_{r}=-\sum_{i}^{n}\sum_{\pi}\log P\left(\bar{S}_{i}\mid N_{i};\theta_{r}\right)$
-After training the schema router, we further design a graph-based constrained decoding algorithm to guarantee the validity of generated schemas.
+Figure 4.
+An example of graph-based constrained decoding.
+The underline (_) denotes the separator between schema elements, and a red cross indicates that the generation of the token is invalid and disallowed.
+We leverage diverse beam search to ensure a more varied set of candidate sequences, represented in different colors in this graph.
+After training the schema router, we further design a graph-based constrained decoding algorithm to guarantee the validity of generated schemata.
 As shown in Figure 4, at each step of autoregressive decoding, we maintain a dynamic prefix tree containing the names of accessible nodes from decoded schema elements.
 Unlike the DFS serialization, the accessible table nodes can only be the neighbors of generated tables, not the neighbors of the database, unless no tables have been generated yet.
-Therefore, the available tokens (part of the node name) for each decoding step can be obtained by searching the prefix tree, prefixed with the tokens generated after the last element separator.
+Therefore, the available tokens (i.e., part of the node name) for each decoding step can be obtained by searching the prefix tree, prefixed with the tokens generated after the last element separator.
 Besides, multiple candidate schema sequences can be generated to improve the coverage of the target schema.
-We thus utilize diverse beam search (Vijayakumar et al., 2016), which allows for generating a more diverse set of candidate sequences.
-Figure 4. An example of graph-based constrained decoding.
+We thus utilize diverse beam search (Vijayakumar et al., 2016) to generate more diverse sets of candidate sequences and combine tables from schema sequences that share the same database to form the final candidate schemata.
 3.6. SQL Generation
-By decoupling schema-agnostic NL2SQL into schema routing and SQL generation, DBCopilot can easily integrate with orthogonal advanced LLM-based NL2SQL solutions, no matter whether in-context prompt engineering methods (Pourreza and Rafiei, 2023; Gao et al., 2023; Guo et al., 2023; Chang and Fosler-Lussier, 2023) or specific NL2SQL LLMs (Defog.ai, 2023; Station, 2023).
-However, since the schema router can generate multiple schemas from different databases for an NL question, we explore three prompt strategies to select and incorporate these different database schemas for LLMs:
-  • Best Schema Prompting: instructs LLMs with the highest probability schema generated by the schema router.
-  • Multiple Schema Prompting: instructs LLMs using multiple candidate schemas by simply concatenating the table schemas from different sequences in the prompt.
-  • Multiple Schema COT Prompting: instructs LLMs using multiple candidate schemas through chain-of-thought (COT) reasoning (Wei et al., 2022).
-The LLM is first asked to select the most relevant candidate schema and then generates SQL queries.
+Basic Prompt ### Complete sqlite SQL query only and with no explanation ### Sqlite SQL tables, with their properties: # # country(code, name, continent, region, ...) # countrylanguage(countrycode, language, ...) # ### Which language is the most popular on the Asian continent?
+SELECT
+Figure 5. An example of a basic NL2SQL prompt.
+By decoupling schema-agnostic NL2SQL into schema routing and SQL generation, DBCopilot can easily integrate with orthogonal, advanced LLM-based NL2SQL solutions, no matter whether in-context prompt engineering methods (Pourreza and Rafiei, 2023; Gao et al., 2024; Guo et al., 2023; Chang and Fosler-Lussier, 2023) or specific NL2SQL LLMs (Defog.ai, 2023; Station, 2023; Li et al., 2024c).
+However, since the schema router can generate multiple schemata from different databases for an NL question, we explore three prompt strategies to select and incorporate these candidate schemata for LLM-based NL2SQL:
+  • Best Schema Prompting: This strategy instructs LLMs with the highest probability schema generated by the schema router, as shown in Figure 5.
+This basic prompt template has been shown to be the most effective prompt for SQL generation with OpenAI LLMs (Gao et al., 2024).
+  • Multiple Schema Prompting: This strategy instructs LLMs with multiple schemata generated by the schema router for more possible coverage of the target schema.
+The prompt format is the same as the best schema prompting strategy and we concatenate multiple table schemata in the prompt.
+  • Multiple Schema COT Prompting: This strategy instructs LLMs using multiple candidate schemata through chain-of-thought (COT) reasoning.
+As shown in Figure 6, the LLM is first asked to select the most relevant candidate schema and then generates SQL queries.
+Chain-of-Thought Prompt Turn 1: Based on the provided natural language question, find the database that can best answer this question from the list schemata below.
+Only output the corresponding database schema identifier in the [id] format, without any additional information.
+Question: Which language is the most popular on the Asian continent?
+Sqlite SQL databases, with their tables and properties: [1] car
+continents(contid, continent)
+countries(countryid, countryname, continent)
+car_makers(id, marker, fullname, country)
+[2] word
+country(code, name, continent, region, ...)
+countrylanguage(countrycode, language, ...)
+[3] ... [4] ... [5] ...
+Turn 2: Filled basic prompt with Turn 1 selected schema.
+Figure 6. An example of a chain-of-thought NL2SQL prompt with multiple candidate schemata.
 4. Experiments
-In this section, we conduct experiments on schema routing and schema-agnostic NL2SQL to verify the effectiveness of DBCopilot.
+In this section, we conduct experiments to verify the feasibility and effectiveness of DBCopilot.
+First, we compare DBCopilot with various retrieval-based methods for schema routing.
+Then, we evaluate the performance of end-to-end schema-agnostic NL2SQL.
+Finally, we perform detailed ablation studies.
 4.1. Experimental Setup
+4.1.1. Research Questions
+We conducted experiments to answer the following key research questions (RQs):
+RQ1: How does DBCopilot compare with various retrieval baselines in schema routing?
+RQ2: Is DBCopilot more robust against semantic mismatch between NL questions and schemata?
+RQ3: Does improving schema routing lead to better end-to-end schema-agnostic NL2SQL?
+RQ4: What is the best prompt strategy for LLM-based SQL generation when there are multiple candidate schemata?
+RQ5: How does each module (i.e., DFS serialization, data synthesis, decoding strategies) contribute to DBCopilot?
 4.1.2. Datasets
 We adapt existing NL2SQL datasets to evaluate schema routing and schema-agnostic NL2SQL.
-  • Spider (Yu et al., 2018): Spider is a widely-used cross-domain NL2SQL dataset with 10,181 questions and 5,693 unique SQL queries across 138 domains and 200 databases.
-  • Bird (Li et al., 2023b): Bird is a recent challenging NL2SQL dataset designed to emphasize the importance of database values, spotlighting the challenges presented by dirty database content, external knowledge requirements, and SQL efficiency.
-  • Fiben (Sen et al., 2020): Fiben is a dataset designed to evaluate the processing of complex business intelligence queries, often involving nesting and aggregation, for NL2SQL systems.
-It consists of 300 NL questions, corresponding to 237 different complex SQL queries on a database with 152 tables.
-  • Spidersyn (Gan et al., 2021): It is a human-curated dataset designed to evaluate the robustness of NL2SQL models to synonym substitution.
-It builds on Spider with NL questions modified by replacing schema-related words with real-world paraphrases.
-  • Spiderreal (Deng et al., 2021): This is another robustness dataset created from the dev split of the Spider dataset by manually modifying the original questions to remove the explicit mention of column names while leaving the SQL queries unchanged.
-Table 2. Statistics of datasets. Asterisks (*) indicate that the original development sets are used as test sets.
-Dataset Size Schema
-Train Dev. Test # DB # Tables # Cols
-Spider 7000 1034* 166 876 4503
-Bird 9427 1534* 80 597 4337
-Fiben - - 279 1 152 374
-Spidersyn 7000 1034 166 876 4503
-Spiderreal 508 166 876 4503
-Dataset Adaptation.
-These datasets are designed to evaluate NL2SQL on a single database, where each question is annotated with its target database.
-To approach the schema-agnostic NL2SQL, we make several adjustments.
-First, we remove the single-database constraint and treat all databases in the collection as potential query targets.
-Next, we extract metadata (tables and columns) from each SQL query using a SQL parser (Mao, 2023) and exclude any queries that cannot be parsed.
-Finally, we combine the target database and extracted metadata as a SQL query schema, together with the original NL question and SQL query, to form each instance ($N$, $S$, $Q$).
-This results in three database collections (Spider, Bird, and Fiben), with instances divided into training and test sets as in the original setup.
-The robustness variants of Spider share the same database collection.
-Table 2 summarizes the dataset statistics after adaptation.
-Table 3. Schema routing performance on standard test sets. The best results are in bold and the second best are underlined.
-Method Spider Bird Fiben
+  • Spider (Yu et al., 2018): Spider is a widely-used, cross-domain NL2SQL dataset with 10,181 questions and 5,693 unique SQL queries across 138 domains and 200 databases, requiring to generalize across different SQL queries and database schemas.
+  • Bird (Li et al., 2023b): Bird is a challenging NL2SQL dataset designed to emphasize the importance of database content, spotlighting the challenges presented by dirty database values, external knowledge requirements, and SQL efficiency.
+  • Fiben (Sen et al., 2020): Fiben is a dataset designed to simulate real-world data mart by IBM that closely mirrors actual enterprise data warehouses.
+Its schema conforms to standard financial ontologies, representing the complexity of schema routing in production environments.
+Fiben consists of 300 NL questions, corresponding to 237 complex SQL queries.
+These queries focus on business intelligence scenarios involving nesting and aggregation, making it an ideal testbed for evaluating NL2SQL systems in real-world scenarios.
+  • Spidersyn (Gan et al., 2021): Spidersyn is a dataset designed to evaluate the robustness of NL2SQL models to synonym substitution, built on top of Spider with NL questions modified by replacing schema-related words with real-world paraphrases.
+  • Spiderreal (Deng et al., 2021): Spiderreal is another robustness dataset created from the dev split of the Spider dataset by manually modifying the original NL questions to remove the explicit mention of column names.
+Table 2. Statistics of datasets.
+{NiceTabular}
+ccccccc Dataset Size Schema
+Train Dev. Test # DBs # Tables # Cols
+[…]
+[…]
+Fiben – – 279 1 152 374
+Spidersyn 7000 – 1034 166 876 4503
+Spiderreal – – 508 166 876 4503
+Table 3. Schema routing performance on regular test sets.
+{NiceTabular}
+[…]
 Database Table Database Table Database Table
@@
 […]
+Dataset Adaptation.
+These datasets are designed to evaluate NL2SQL on a single database, where each question is annotated with its target database.
+To approach the schema-agnostic NL2SQL, we make several adjustments.
+First, we remove the single-database constraint and treat all databases in the collection as potential query targets.
+Next, we extract metadata (tables and columns) from each SQL query using a SQL parser (Mao, 2023) and exclude any queries that cannot be parsed.
+Finally, we combine the target database and extracted metadata as a SQL query schema, along with the original NL question and SQL query, to form each instance ($N$, $S$, $Q$).
+This results in three database collections (Spider, Bird, and Fiben), with instances divided into training and test sets as in the original setup.
+The robustness variants of Spider share the same database collection.
+§ 4.1.2 summarizes the dataset statistics after adaptation.
 4.1.3. Baselines
-We compare DBCopilot to various sparse/dense retrieval techniques for scaling NL querying over massive databases, including zero-shot, LLM-enhanced, and fine-tuned methods:
-  • BM25 (Wikipedia contributors, 2023a): BM25 is a well-known ranking function widely used in search systems to retrieve relevant documents.
+We compare DBCopilot to various sparse and dense retrieval techniques for schema routing (RQ1 & RQ2), including zero-shot, LLM-enhanced, and fine-tuned methods:
+  • BM25 (Wikipedia contributors, 2023b): BM25 is a renowned ranking function that serves as a standard in search engines to rank documents by their relevance to a specified query.
   • SXFMR (Reimers and Gurevych, 2019): SXFMR is a technique for obtaining generic embedding models for dense retrieval by contrastive learning of Transformer(XFMR)-based PLMs on related text pairs.
   • CRUSH (Kothyari et al., 2023): CRUSH adopts a two-stage approach to retrieve schema elements for NL2SQL.
-It first instructs the LLM to generate a hallucinated schema, and then combines multiple retrieval results to obtain the actual schema.
-  • DTR (Herzig et al., 2021): DTR is an approach that uses contrastive learning on labeled (question, table) pairs to develop table retrievers for open-domain question answering over tables.
+It first instructs an LLM to generate a hallucinated schema, and then combines multiple retrieval results and reranks them based on their relationships to obtain the actual schema.
+  • DTR (Herzig et al., 2021): DTR is an approach that uses contrastive learning on (question, table) pairs to develop table retrievers for open-domain question answering over tables.
+For schema-agnostic NL2SQL (RQ3 & RQ4), since existing NL2SQL methods cannot be directly applied to this problem even when equipped with schema linking (as discussed in § 2.1), we combine a representative LLM-based NL2SQL method (Gao et al., 2024) with different retrieval methods for comparison.
+This allows us to isolate and evaluate the effects of schema routing and prompt strategies on end-to-end SQL generation in a controlled setting.
 4.1.4. Evaluation Metrics
+Different evaluation metrics are used for schema routing and SQL generation respectively.
 For schema routing, we use Recall@k and mAP (mean average precision) to evaluate the relevance of retrieved databases and tables.
@@
 For SQL generation, traditional surface-level metrics are inappropriate for LLM-based approaches as they may produce SQL queries that differ greatly from the ground truth.
-Following recent studies (Gu et al., 2023b; Gao et al., 2023; Li et al., 2023b), we use EX (execution accuracy), which compares the results of generated SQL queries with corresponding ground truths.
-We also report the cost of LLM invocations.
+Following recent studies (Gu et al., 2023b; Gao et al., 2024; Li et al., 2023b), we use EX (execution accuracy), which compares the results of generated SQL queries with corresponding ground truths.
+We also report the cost (\$) of LLM invocations.
 4.1.5. Implementation Details
-We implemented DBCopilot using Python 3.10, Pytorch 2.1, and the Transformers (Wolf et al., 2020) library.
+We implemented DBCopilot using Python 3.10, Pytorch 2.1, along with some major libraries such as Transformers (Wolf et al., 2020) and Pytorch Lightning (Falcon and The PyTorch Lightning team, 2019).
+The source code is available at github.com/tshu-w/DBCopilot.
 Schema Routing.
 We adopted T5-base (Raffel et al., 2020) as the backbone of the schema questioner and router.
+We trained a unified schema questioner on Spider and Bird NL2SQL training sets and three schema routers on each database collection (Spider, Bird, and Fiben).
+For each database collection, we sampled $1\times 10^{5}$ SQL query schemata uniformly from the constructed graph, covering all (100%) databases and tables, and synthesized the corresponding training instances using the schema questioner.
 We set the batch size to $32$, the learning rate to $5\times 10^{-5}$, and the maximum number of training epochs to $20$ for all models.
 Model parameters were optimized using AdamW (Loshchilov and Hutter, 2019) and a linear learning rate schedule with no warm-up steps.
-We used the schema questioner trained on Spider and Bird to generate pseudo-questions for Fiben, which has no training data.
-For each database collection, we synthesized $10^{5}$ training instances.
+To detect joinable tables besides primary and foreign, we adopted a heuristic way that two tables are joinable if the exact match overlap (Jaccard similarity) of their column values is greater than 0.85.
+During schema routing, we generated 10 schema sequences for each question using diverse beam search, with beams and beam groups set to 10 and the diversity penalty set to 2.0.
 SQL Generation.
-We used gpt-3.5-turbo-16k-0613 to select the optimal schema and generated the SQL query with the OpenAI demo prompt (Gao et al., 2023).
-We set the temperature to 0 for reproducibility.
+We used gpt-3.5-turbo-0125222https://platform.openai.com/docs/models#gpt-3-5-turbo to select the optimal schema and generated the SQL query, with the generation temperature set to 0 for reproducibility.
 Baselines.
@@
 We treated tables as retrieval targets, with the content of each table being the flat normalized names of the table and its columns.
-For each NL question, we retrieved the top 100 tables and determined database scores based on the average of their table scores.
+For each NL question, we retrieved the top 100 tables and ranked the databases based on the average score of their retrieved tables.
 We used the standard Okapi BM25 with two adjustable parameters for BM25 and adopted the best performing all-mpnet-base-v2 for SXFMR.
-BM25 and DTR were fine-tuned on synthetic data, consistent with DBCopilot.
-4.2. Schema Routing
-We first compare the performance of schema routing, with the results shown in Table 3, Table 4, Figure 5, and the following findings.
-Table 3 shows the database and table recall on regular test sets.
-We can see that DBCopilot consistently outperforms all types of baselines by a large margin.
+BM25 and DTR were fine-tuned on synthetic data consistent with DBCopilot to verify our relation-aware joint retrieval.
+All experiments were run using Docker image nvidia/cuda:12.1.1-cudnn8-devel-ubuntu22.04 on a server with AMD EPYC 7H12 64-Core Processors, 1TB RAM, and NVIDIA A100 40G GPUs.
+4.2. Schema Routing (RQ1 & RQ2)
+4.2.1. Experimental Results
+We first compare the performance of schema routing on both regular and robustness datasets, with the experimental results shown in § 4.1.2, § 4.2.1, Figure 7, and the following findings.
+§ 4.1.2 shows the database and table recall on regular test sets.
+We can see that DBCopilot consistently outperforms all zero-shot, LLM-enhanced, or fine-tuned baselines by a large margin.
 The database recall@1 of DBCopilot surpasses the second-best approach by up to 12.67% and 11.93% on the Spider and Bird datasets, respectively.
@@
 Finding 1. DBCopilot consistently outperforms all types of baselines by a large margin in schema routing.
-Table 4. Schema routing performance on robustness test sets.
-Method Spidersyn Spiderreal
+When querying over massive databases, data consumers without schema knowledge typically express NL questions that are highly inconsistent with database schemata.
+To evaluate the robustness of different methods in this situation, § 4.2.1 shows the experimental results on Spidersyn and Spiderreal, which replace explicitly mentioned schema elements in NL questions posed by NL2SQL annotation.
+We can see that the performance of all methods decreases in this actual scenario setting.
+Sparse retrieval baselines are significantly affected as they rely heavily on the vocabulary matching, and the LLM-enhanced method also struggles with this challenge, despite being designed specifically for this case.
+Compared to these baselines, DBCopilot is the least affected and therefore maintains a larger advantage with 14.89% and 19.88% on database recall@1, as well as 2.83% and 6.94% on table recall@5.
+We believe that by injecting schema knowledge through schema-specific training, DBCopilot can better understand the semantics of database schemata and effectively learn the semantic mapping between NL questions and schemata.
+Table 4. Schema routing performance on robustness tests.
+{NiceTabular}
+[tabularnote = We use the same indicators and markups as § 4.1.2.]@ccccccccc@ Method Spidersyn Spiderreal
 Database Table Database Table
@@
 […]
-DBCopilot 62.67 85.11 70.35 86.26 79.53 95.47 87.93 96.06
-When querying over massive databases, data consumers without schema knowledge typically express NL questions that are highly inconsistent with database schemas.
-To evaluate the robustness in this situation, Table 4 shows the results of Spidersyn and Spiderreal, which replace explicitly mentioned schema elements in NL questions posed by NL2SQL annotation.
-We can see that the performance of all methods decreases with this diversity setting.
-Sparse retrieval baselines are significantly affected as they rely heavily on the vocabulary matching assumption, and the LLM-enhanced method also struggles with this challenge, despite being designed specifically for this case.
-Compared to these baselines, DBCopilot is the least affected and therefore maintains a larger advantage with 14.89% and 19.88% on database recall@1, as well as 2.83% and 6.94% on table recall@5.
-We believe this is because by injecting schema knowledge through schema-specific training, our method can better understand the semantics of database schemas and effectively learn the mapping between questions and schemas.
-Finding 2. DBCopilot is more robust when natural language questions are highly inconsistent with database schemas.
-Figure 5. (a) Table mAP w.r.t. different numbers of database tables. (b) Table recall@$k$ w.r.t. the number of retrieved tables.
-DBCopilot model schema routing as a relation-aware, generative joint retrieval process, which differs from traditional baselines of independent retrieval.
-To assess the effect of relation modeling and joint retrieval, we evaluate the performance under different numbers of database tables and retrieved tables.
-As shown in Figure 5 (a), the mAP of DBCopilot is relatively stable as the number of database tables increases, while the independent retrieval baselines decrease significantly.
-The recall of DBCopilot is also consistently higher than the baselines as the number of retrieved tables increases from Figure 5 (b).
-These results verify the scalability of DBCopilot through relation modeling and generative routing.
-Finding 3. Relation modeling and joint retrieval are essential for DBCopilot to perform effective schema routing.
-4.3. Schema-Agnostic NL2SQL
-We then compare the schema-agnostic NL2SQL performance under gold schemas, different routing methods, and prompt strategies.
+DBCopilot 62.67♯‡ 85.11 70.35♯‡ 86.26♯‡ 79.53♯‡ 95.47♯‡ 87.93♯‡ 96.06♯‡
+Finding 2. DBCopilot is more robust when NL questions are highly inconsistent with database schemata.
+Figure 7. (a) Table mAP w.r.t. different numbers of database tables. (b) Table recall@k w.r.t. different numbers of retrieved tables.
+We further analyzed the performance regarding different database sizes (i.e., the number of tables contained) and the number of tables retrieved.
+Experimental results are shown in Figure 7 (a) and (b), respectively.
+First, we categorize the instances based on the number of tables contained in their target database and compare the table mAP across different database sizes.
+As shown in the Figure 7 (a), the mAP of DBCopilot is relatively stable as the number of database tables increases, while the independent retrieval baselines decrease significantly.
+We believe this is because by modeling schema routing as a relation-aware joint retrieval process, DBCopilot can better capture the relationships within the database compared to independent retrieval methods, which is critical for routing databases that contain many tables.
+Furthermore, as shown in Figure 7 (b), the table recall of DBCopilot is consistently higher than that of the baselines as the number of retrieved tables increases, achieving a recall@20 that exceeds the recall@50 of all baselines.
+This suggests that retrieving more tables and reranking them based on their relationships can hardly compensate for the advantages of our end-to-end relation-aware joint retrieval approach.
+Finding 3. Relation-aware joint retrieval are essential for DBCopilot to perform effective schema routing.
+4.2.2. Case Study
+[…]
+Figure 8. Successful case of DBCopilot in schema routing.
+We provide successful cases to illustrate how DBCopilot improves schema routing through schema-specific training and joint retrieval.
+As shown in Figure 8, DBCopilot is the only approach that accurately routes the appropriate schema for the given question.
+Sparse retrieval methods (i.e., BM25 and CRUSHBM25) rely heavily on lexical matching while ignoring the semantic matching between the question and the schemata.
+However, untuned dense retrieval methods (i.e., SXFMR and CRUSHSXFMR) also fail to capture the underlying semantics of tables and retrieve the wrong databases.
+Fine-tuning based approaches identify the correct database where DTR misses the country table linked to the countrylanguage table, but our relation-aware DBCopilot does not.
+While DBCopilot consistently outperforms the baselines, there are a few cases where DTR retrieves the correct schema but DBCopilot does not.
+For example, Figure 9 shows that DBCopilot confuses the airlines and airports tables of database flight, but DTR covers the correct tables by retrieving more loosely.
+Upon further analysis, we find that the quality of synthetic data impacts the robustness of schema routing.
+This is attributed to two main factors: 1) hallucination (Papicchio et al., 2023), where the generated NL questions deviate from sampled schemata; 2) generation bias, where the simple pipeline lacks diversity and quality control mechanisms.
+Developing more sophisticated data synthesis techniques to address these challenges remains an important direction for future work.
+[…]
+[…]
+[…]
+[…]
+Table 5. Method efficiency & resource consumption.
+[…]
+[…]
+BM25 1677.1 10 36 800 -
+SXFMR 49.0 48 420 2,230 3742
+[…]
+[…]
+BM25 1677.1 6972 36 800 -
+DTR 49.0 5161 420 2,230 3742
+DBCopilot 43.0 13678 852 656 11807
+4.2.3. Efficiency & Resource Consumption
+We also measure the efficiency and resource consumption of each schema routing method.
+§ 4.2.2 shows the number of queries per second (QPS), the duration of build time, the amount of disk space used, and the consumption of system and GPU memory.
+LLM inference and complex post-processing are the main reasons for the slowness of CRUSH approaches.
+DBCopilot, based on an encoder-decoder architecture, consumes more disk space and GPU memory than dense retrieval approaches that based on an encoder-only architecture.
+Despite the slightly higher resource consumption for training and inference in generative retrieval, we believe and have observed that the growing popularity of generative AI is leading to an increasing number of proposed optimizations.
+4.3. Schema-Agnostic NL2SQL (RQ3 & RQ4)
+This section compares the performance of schema-agnostic NL2SQL under various situations, including providing gold schemata, using different schema routing methods, and prompt strategies.
+Experimental results are shown in § 4.3.
 Oracle Test.
-We conduct oracle tests to access the upper bound of schema-agnostic NL2SQL, by providing gold schemas at various small scales.
-Specifically, we gradually move from 5 database schemas to the gold database, then to gold tables only, or even to gold tables with only gold columns.
-Experimental results in Table 5 show that as the number of schema elements increases, the EX of LLM-generated SQL queries continuously decreases, while the inference cost keeps increasing.
+We first conduct oracle tests to access the upper bound of schema-agnostic NL2SQL, as well as the impact of irrelevant schema elements for LLM-based NL2SQL, by providing gold schemata at various small scales.
+Specifically, we gradually move from 5 database schemata to the gold database, then to gold tables only, or even to gold tables with only gold columns.
+Experimental results in § 4.3 show that as the number of schema elements increases, the EX of LLM-generated SQL queries continues to decrease significantly, while the inference cost keeps increasing many times over.
 This suggests that LLMs, like PLMs (Li et al., 2023c), are also affected by extraneous schema elements in SQL generation.
-Predictably, the situation gets worse as the scale of the schema increases.
-As a result, it is highly inefficient and ineffective to expose massive database schemas directly to LLMs.
-Table 5. Schema-agnostic NL2SQL performance with gold (G.) schemas, different routing methods, and prompt strategies.
-Routing Method Provided Schema Prompt Strategy Spider Bird Spidersyn
+Predictably, the situation gets worse as the schema scales.
+As a result, it is highly inefficient and ineffective to expose massive database schemata directly to LLMs, making schema routing necessary for schema-agnostic NL2SQL.
+Table 6. Evaluation of schema-agnostic NL2SQL with various gold schemata, different schema routing methods, and prompt strategies.
+[…]
+notes/style=0 {NiceTabular}[tabularnote = “T.” and “C.” stand for table and column, respectively.
+We bold best results of schema-agnostic NL2SQL and underline the best with human in the loop to select the optimal candidate schema.
+For retrieval baselines, a candidate schema consists of the top candidate database and the retrieved tables that belong to the database.
+]@cccccccc@ Routing Method Prompt Template Spider Bird Spidersyn
 EX Cost EX Cost EX Cost
-G. Column Basic 79.69 0.40 35.98 0.75 76.60 0.40
-G. Table Basic 77.08 0.52 29.86 1.44 67.21 0.52
-G. Database Basic 74.47 0.81 25.75 2.70 63.73 0.81
-5 Database Basic 67.89 2.98 19.10 8.80 49.90 2.99
-CRUSHBM25 Top 1 Basic 56.77 1.30 13.69 2.67 20.79 1.23
-DTR Top 1 Basic 45.16 0.66 17.60 2.12 26.98 0.62
-DBCopilot Top 1 Basic 64.12 0.69 21.97 2.35 37.81 0.70
-DBCopilot Top 5 Basic 61.51 2.28 19.56 7.18 34.53 2.27
-DBCopilot Top 5 COT 57.83 5.33 19.17 16.49 33.95 5.29
+Oracle Test
+Gold T. & C. 81.24 0.07 35.98 0.15 77.37 0.07
+Gold T. 76.60 0.09 31.03 0.24 68.67 0.09
+Gold DB 74.08 0.12 27.57 0.40 61.99 0.13
+5 DB w. Gold 70.21 0.13 18.38 8.80 51.64 0.40
+Best Schema Prompting
+CRUSHBM25 Top 1 56.58 0.10 14.73 0.26 20.60 0.09
+DTR Top 1 43.91 0.11 19.04 0.33 26.21 0.10
+DBCopilot Top 1 63.35 0.11 23.47 0.36 37.43 0.11
+Multiple Schema Prompting
+CRUSHBM25 Top 5 55.61 0.20 17.60 0.57 21.86 0.19
+DTR Top 5 54.45 0.27 19.56 0.79 32.59 0.27
+DBCopilot Top 5 61.51 0.31 21.45 0.95 35.49 0.31
+Multiple Schema COT Prompting
+CRUSHBM25 Top 5 53.97 0.28 22.10 0.83 22.34 0.26
+DTR Top 5 50.68 0.36 22.43 1.08 28.53 0.37
+DBCopilot Top 5 51.26 0.40 22.75 1.24 31.24 0.39
+Human in the Loop
+CRUSHBM25 Top 5 69.34 0.10 25.10 0.32 34.72 0.09
+DTR Top 5 66.92 0.11 25.42 0.37 47.78 0.11
 […]
-Finding 4. Schema routing is necessary for LLM-based schema-agnostic NL2SQL for both efficiency and effectiveness.
+Finding 4. Schema routing is necessary for scaling natural language querying to massive databases.
 Impact of Improved Recall.
-We perform schema-agnostic NL2SQL to verify the impact of improved recall on the accuracy of generated SQL queries.
-Without loss of generality, we compare DBCopilot to CRUSHBM25 (sparse) and DTR (dense), which achieve suboptimal routing performance on Spider and Bird, respectively.
-As shown in Table 5, DBCopilot outperforms these two baselines by up to 10.83% EX.
-The LLM inference cost of DBCopilot is slightly higher, possibly due to successfully retrieving more required tables.
-We can also confirm the correlation between routing recall and EX by comparing these two baselines.
+We then perform schema-agnostic NL2SQL to verify the impact of improved recall on the accuracy of generated SQL queries.
+Without loss of generality, we compare DBCopilot to sparse CRUSHBM25 and dense DTR baselines, which achieve suboptimal routing performance on Spider and Bird, respectively.
+As shown in § 4.3, DBCopilot outperforms CRUSHBM25 and DTR by 4.43%~11.22% EX when using the best schema prompting strategy and by 0.85%~20.07% EX when employing human in the loop to select the optimal schema from candidates.
+We compare and discuss the prompting strategies in the next paragraph.
+The LLM inference cost of DBCopilot is slightly higher, possibly due to its successful retrieval of more required tables.
+The correlation between schema routing recall and SQL generation EX can also be confirmed by comparing these two baselines.
 CRUSHBM25 outperforms DTR in recall and EX on Spider, while DTR achieves higher EX on the other two datasets due to its superior recall.
-Therefore, the success of schema routing is essential for schema-agnostic NL2SQL.
-Finding 5. DBCopilot outperforms retrieval baselines in schema-agnostic NL2SQL due to its superior schema routing.
+Thus, the success of schema routing is essential for schema-agnostic NL2SQL.
+Finding 5. DBCopilot leads in schema-agnostic NL2SQL due to its superior schema routing performance.
 Exploration of Prompt Strategies.
-By generating multiple candidate schemas, DBCopilot can effectively increase the recall of the target schema.
-We explore three prompt strategies for leveraging multiple schemas.
-Interestingly, SQL generation with the top 1 schema performs best, even though the top 5 candidates have 22.44% more database recall on Spidersyn than the top 1.
-This may be due to confusion caused by similar candidate schemas, thus requiring more database/human information to disambiguate.
-By introducing human-in-the-loop (HITL) to select the desired schema from the top 5 candidates, DBCopilot comes close to the ideal situation of providing the gold database schema, but with lower inference costs.
-This highlights the ability of DBCopilot to exclude extraneous tables and its promise for schema-agnostic NL2SQL.
-Finding 6. Using the best schema prompting strategy for SQL generation is optimal unless there is human interaction.
-4.4. Ablation Study
-Table 6. Performance loss in ablation studies. “BS”, “OD”, “CD”, “DB” stands for basic serialization, original NL2SQL training data, constrained decoding, and diverse beam search.
-Spider Bird
+By routing multiple candidate schemata, the recall of the target schemata for NL questions can be greatly improved.
+We explore three prompt strategies for leveraging multiple candidate schemata during SQL generation.
+As shown in § 4.3, when schema routing through DBCopilot, the performance of multiple schema prompting is slightly lower than that of single schema prompting.
+However, when schema routing through baselines, the performance of multiple schema prompting is typically better than that of single schema prompting.
+This is due to the trade-off between the improved recall brought by multiple candidate schemata and the impact of extraneous schemata on LLMs.
+As for multiple schema COT prompting, this strategy shows no absolute advantage over multiple schema prompting, and each wins and loses on different routing methods and datasets.
+Thus, the limitations of LLMs in distinguishing domain-specific schemata make it cost effective for DBCopilot to find the correct schema in the top 1 candidate.
+Finally, by introducing human-in-the-loop to select the desired schema from top 5 candidates, DBCopilot comes close to the ideal performance of providing the gold database schema, but with lower costs.
+Finding 6. The best schema prompting strategy is optimal for DBCopilot unless there is human interaction.
+4.4. Ablation Study (RQ5)
+Table 7. Performance loss in ablation studies.
+[…]
+notes/style=0 {NiceTabular}@lcccccccc@ Spider Bird
 Database Table Database Table
@@
 DBCopilot 85.01 96.42 91.63 97.51 88.92 97.13 85.83 94.59
-   w/ BS -3.20 -2.90 -4.09 -5.75 -3.39 -1.63 -6.43 -5.99
-   w/ OD -84.72 -82.40 -88.76 -79.55 -88.92 -82.27 -85.73 -88.81
-   w/o CD -0.39 -0.19 -4.15 -2.51 +0.12 -0.45 -2.81 -1.93
-   w/o DB +0.48 -3.00 +0.64 -2.88 -0.52 -4.89 -3.69 -8.87
-We finally perform ablation studies on DBCopilot to validate the contribution of each module, as summarized in Table 6.
-DFS Serialization.
-DFS serialization is essential for schema routers to learn the underlying relations of databases and tables.
-To demonstrate this, we replace the DFS serialization with the unordered basic serialization of tables, similar to DSI for multiple documents.
-As shown in Table 6, this results in a 1% to 6% reduction in recall.
-DFS serialization effectively reduces the search space, whether by first generating a database and then generating its tables, or by continuing to generate based on existing tables.
-This graph-based serialization also makes constrained decoding possible.
-Data Synthesis.
-Generative retrieval is difficult to generalize to unseen documents during training.
-As we can see, routers trained on the original training instances ($N$, $S$) extracted from NL2SQL datasets are almost completely incapable of routing on test schemas, even when constrained decoding is used.
-Data synthesis by sampling schemas and generating pseudo-questions can effectively solve this challenge.
-Figure 6 shows how the performance varies with the amount of synthetic data.
-Database and table recalls increase rapidly at first as more data is synthesized, but level off after a certain amount.
-How to improve the quality of synthetic data and reduce the need for it may be a future direction.
-Decoding Strategy.
-Decoding optimizations (i.e., constrained decoding and group beam search) are beneficial for generative schema routing during inference.
-On the one hand, constrained decoding is useful for generating accurate table identifiers, since the drop from removing it is mainly in table recalls.
-On the other hand, group beam search helps to generate multiple diverse schemas because database recall@5 decreases significantly without it.
-Finding 7. DFS serialization, data synthesis, and decoding optimizations contribute to DBCopilot in different ways.
-5. Related Work
-5.1. Natural Language to SQL
-NL2SQL is an essential part of the natural language interface.
-Most research has approached this task along the lines of semantic parsing (Kamath and Das, 2019; Li et al., 2020).
-Early NL2SQL parsing systems adopted symbolic approaches that relied heavily on rule or grammar engineering, making them difficult to apply widely (Zelle and Mooney, 1996; Woods, 1973).
-In recent years, with the advances in deep learning, especially PLMs, end-to-end NL2SQL approaches that using neural generation models have become the mainstream (Li et al., 2023c; Wang et al., 2020; Gu et al., 2023a; Li et al., 2023a; Yu et al., 2021).
-Some efforts have been made to improve encoding and decoding, such as relation-aware encoding (Wang et al., 2020; Li et al., 2023a) and grammar-based decoding (Xu et al., 2017; Scholak et al., 2021).
-Some other efforts have focused on generating accurate SQL queries through schema linking (Li et al., 2023c; Wang et al., 2020) or skeleton generation (Gu et al., 2023a, b).
-LLMs have emerged as a new paradigm for NL2SQL (Rajkumar et al., 2022; Liu et al., 2023a; Gao et al., 2023).
-Unlike previous work, the core of the LLM-based NL2SQL solution is prompt engineering, i.e., prompting LLM to generate correct and efficient SQL queries.
-Several efforts have been made to select appropriate in-context learning examples through question classification (Pourreza and Rafiei, 2023), skeleton retrieval (Guo et al., 2023), and in-domain example synthesis (Chang and Fosler-Lussier, 2023).
-Research has also been done on decoupling zero-shot NL2SQL into PLM-based sketch generation and LLM-based SQL completion (Gu et al., 2023b).
-These studies are orthogonal to our work and can be easily integrated to optimize DBCopilot SQL generation.
-5.2. Table Discovery
-Table discovery techniques are critical to data management, allowing users to explore and gain insights from data warehouses or lakes (Fan et al., 2023a).
-A common practice for table discovery is query-driven discovery (Zhang and Ives, 2020; Nargesian et al., 2019), where the query is typically in the form of keywords, tables, or NL questions.
-For keyword search, data discovery systems find relevant tables by searching for topic terms in the metadata (Cafarella et al., 2009; Pimplikar and Sarawagi, 2012).
-For table search, tables are retrieved to augment a given table as the query (Sarma et al., 2012).
-Joinable table search (Dong et al., 2021; Zhu et al., 2019) and unionable table search (Fan et al., 2023b; Khatiwada et al., 2023) aim to augment a query table with additional attributes and new tuples, respectively.
-For NL question search, open domain question answering retrieves independent tables to answer the question with table-based models (Chen et al., 2021; Herzig et al., 2021).
-Although dense retrieval methods based on contrastive learning have achieved advanced performance in the above tasks (Dong et al., 2021; Fan et al., 2023b; Herzig et al., 2021), they are difficult to consider the structural relationships between tables and retrieve them jointly.
-Alternatively, data navigation allows users to navigate interactively through the tables organized hierarchically (Nargesian et al., 2020) or relationally (Fernandez et al., 2018).
-Our work can be seen as automated data navigation, where we employ DSI to model the query-based navigation in an end-to-end manner.
-Figure 6. Database recall@1 and table recall@5 w.r.t. the number of (#) synthetic training data.
-6. Conclusion
-In this paper, we present DBCopilot, a framework for effectively scaling natural language querying to massive databases.
-DBCopilot decouples schema-agnostic NL2SQL into schema routing and SQL generation through LLM-Copilot collaboration.
-We present an approach to effectively navigate the target database and tables for an NL question in a relation-aware, end-to-end manner.
-We also propose a training data synthesis paradigm to adapt routers over massive databases without manual intervention, and explore various prompt strategies to incorporate multiple candidate schemas.
-Extensive experiments verify the effectiveness of our solution.
-In summary, DBCopilot pushes the boundaries of NL2SQL, enabling researchers to better strategize their quest for data accessibility.
-Appendix A Supplementary Experiments
-A.1. Case Study
 […]
@@
 […]
-DPR $\langle\text{world},\{\text{countrylanguage}\}\rangle$
-[…]
-Figure 7. Successful case where DBCopilot routes the correct schema but baselines fail.
-[…]
-[…]
-[…]
-[…]
-We provide successful cases to illustrate how DBCopilot improves schema routing through joint retrieval.
-As the example in Figure 7 shows, DBCopilot is the only approach that accurately routes the appropriate schema for the given natural language question.
-Sparse retrieval methods (BM25 and CRUSHBM25) rely heavily on lexical matching while ignoring the semantic matching between the question and the schemas.
-However, untrained dense retrieval methods (SXFMR and CRUSHSXFMR) also fail to capture the underlying semantics of tables and retrieve the wrong databases.
-Fine-tuning based approaches identify the correct database where DPR misses the country table linked to the countrylanguage table, but relation-aware DBCopilot does not.
-[…]
-A.2. Efficiency & Resource Consumption
-We also measure the efficiency and resource consumption of each schema routing method.
-Table 7 shows the number of queries per second (QPS), the duration of build time, the amount of disk space utilized, and the consumption of system and GPU memory.
-LLM inference and complex post-processing are the main reasons for the slowness of CRUSH approaches.
-DBCopilot, based on an encoder-decoder architecture, consumes more disk space and GPU memory than dense retrieval approaches (SXFMR, DPR) based on an encoder-only architecture.
-Despite the slightly higher resource consumption for training and inference in generative retrieval, we believe and have observed that the growing popularity of generative AI is leading to an increasing number of optimizations being proposed.
-Table 7.
-Efficiency & resource consumption of different methods.
-Build time includes the time for model training and index construction.
-We calculate queries per second (QPS) and memory consumption with a query batch size of 64.
-Method QPS Build (s) Disk (MB) RAM (MB) GPU (MB)
-BM25 1677.1 10 36 800 -
-SXFMR 49.0 48 420 2,230 3742
-CRUSHBM25 0.9 10 36 1104 -
-CRUSHSXFMR 0.8 48 420 2502 3742
-BM25 1677.1 6972 36 800 -
-DPR 49.0 5161 420 2,230 3742
-DBCopilot 43.0 13678 852 656 11807
-A.3. Oracle Test with In-Context Learning
-Table 8. Schema-Aware NL2SQL with In-Context Learning
-Provided Schema Prompt Strategy Spider Bird Spidersyn
-EX Cost EX Cost EX Cost
-G. Column Basic 77.95 3.52 35.98 11.58 76.69 3.43
-G. Table Basic 75.82 3.61 28.49 12.10 69.34 3.52
-G. Database Basic 75.73 3.83 25.75 13.08 63.64 3.74
-5 Database Basic 68.38 5.57 21.45 17.79 50.48 5.47
-Table 8 shows the results of the oracle test by providing few-shot (5-shot) examples.
-As we can see, while this improves the accuracy of NL2SQL somewhat, it does not compensate for the performance degradation caused by extraneous schema elements.
-This makes schema linking still important in LLM-based NL2SQL.
-In the future, we will investigate how to better disambiguate and link schema elements in the context of schema-agnostic NL questions.
+We perform ablation studies on DBCopilot to validate the contribution of each module, as summarized in § 4.4.
+DFS Serialization.
+DFS serialization is essential for schema routers to learn the underlying relations of databases and tables.
+To demonstrate this, we replace the DFS serialization with the unordered basic serialization of tables, similar to DSI for multiple documents.
+As shown in § 4.4, this results in a 1% to 6% reduction in recall.
+In addition, DFS serialization facilitates graph-based constrained decoding to reduce the search space, either by first generating a database and then generating its tables, or by further generating based on existing tables.
+Data Synthesis.
+Generative retrieval enables relation-aware end-to-end schema routing, but is unable to generalize to unseen schemata during training.
+We can see that routers trained on instances ($N$, $S$) extracted from original NL2SQL datasets are almost completely incapable of routing on unseen test schemata.
+Data synthesis by sampling schemata and generating pseudo-questions can effectively inject semantic mappings to model parameters.
+Besides, mixing original and synthetic data for training does not lead to more accurate schema routing, possibly due to label imbalance.
+Figure 10 shows how the performance varies with the amount of synthetic data.
+Database and table recalls increase rapidly at first as more data is synthesized, but level off after a certain amount.
+How to improve the quality of synthetic data and reduce the need for its quantity may be a future direction.
+Decoding Strategy.
+Decoding optimizations (i.e., constrained decoding and group beam search) are beneficial for generative schema routing during inference.
+Specifically, constrained decoding is useful for generating accurate table identifiers, since removing it mainly affects table recalls, about 2% to 4%.
+Meanwhile, group beam search helps to generate multiple diverse schemata because database recall@5 decreases significantly without it.
+Finding 7. DFS serialization, data synthesis, and decoding optimizations contribute in different ways.
+Figure 10. Database recall@1 and table recall@5 w.r.t. the number of (#) synthetic training data.
+5. Related Work
+5.1. Natural Language to SQL
+NL2SQL is an essential part of natural language interfaces to databases, enabling non-experts to easily query structured data.
+Most research has approached this task along the lines of semantic parsing (Kamath and Das, 2019; Li et al., 2020).
+Early NL2SQL parsing systems adopted symbolic approaches that relied heavily on rule or grammar engineering, making them difficult to apply widely (Zelle and Mooney, 1996; Woods, 1973).
+In recent years, with the advances in deep learning, especially PLMs, NL2SQL approaches that using neural generation models have become the mainstream (Li et al., 2023c; Wang et al., 2020; Gu et al., 2023a; Li et al., 2023a; Yu et al., 2021).
+Some efforts have been made to improve encoding and decoding, such as relation-aware encoding (Wang et al., 2020; Li et al., 2023a) and grammar-based decoding (Xu et al., 2017; Scholak et al., 2021).
+Some other efforts have focused on generating accurate SQL queries through schema linking (Li et al., 2023c; Wang et al., 2020) or skeleton generation (Gu et al., 2023a; Gu et al., 2023b).
+LLMs have emerged as a new paradigm for NL2SQL (Rajkumar et al., 2022; Liu et al., 2023; Gao et al., 2024; Li et al., 2024a).
+Unlike previous work, a core component of the LLM-based NL2SQL solution is prompt engineering, i.e., prompting LLM to generate correct and efficient SQL queries.
+Several efforts have been made to select appropriate in-context learning examples through question classification (Pourreza and Rafiei, 2023), skeleton retrieval (Guo et al., 2023), and in-domain example synthesis (Chang and Fosler-Lussier, 2023).
+Research has also been done on decoupling zero-shot NL2SQL into PLM-based sketch generation and LLM-based SQL completion (Gu et al., 2023b).
+Recently, pipeline-based approaches decompose NL2SQL into steps including schema linking, query classification, query decomposition, SQL refinement, further extending the power of LLM-based NL2SQL (Pourreza and Rafiei, 2023; Li et al., 2024b).
+These studies are orthogonal to our work and can be easily integrated to optimize SQL generation of DBCopilot.
+5.2. Table Discovery
+Table discovery techniques are critical to data management, allowing users to explore and gain insights from data warehouses or lakes (Fan et al., 2023a).
+A common practice for table discovery is query-driven discovery (Zhang and Ives, 2020; Nargesian et al., 2019), where the query is typically in the form of keywords, tables, or NL questions.
+For keyword search, data discovery systems find relevant tables by searching for topic terms in the metadata (Cafarella et al., 2009; Pimplikar and Sarawagi, 2012).
+Table understanding (Deng et al., 2020), including domain discovery (Li et al., 2017; Ota et al., 2020) and table annotation (Limaye et al., 2010; Venetis et al., 2011), has emerged as a means to mitigate the missing, incomplete, and inconsistent table schemata for keyword search.
+For table search, tables are retrieved to augment a given table as the query (Sarma et al., 2012).
+Joinable table search (Dong et al., 2021; Zhu et al., 2019) and unionable table search (Fan et al., 2023b; Khatiwada et al., 2023) aim to augment a query table with additional attributes and new tuples, respectively.
+For NL question search, open domain question answering retrieves independent tables to answer the NL question with table-based models (Chen et al., 2021; Herzig et al., 2021).
+Although dense retrieval methods based on contrastive learning have achieved advanced performance in the above tasks (Dong et al., 2021; Fan et al., 2023b; Herzig et al., 2021), they are difficult to consider the structural relationships between tables and retrieve them jointly.
+Alternatively, data navigation allows users to navigate interactively through the tables organized hierarchically (Nargesian et al., 2020) or relationally (Fernandez et al., 2018).
+Our work can be seen as automated data navigation, where we employ relation-aware DSI to model the query-based navigation in an end-to-end manner.
+There is another line of research that focuses on table content discovery for Web table QA (Badaro et al., 2023), which typically employs PLMs to manipulate and analyze HTML tables of limited size (with average #row < 20) (Hulsebos et al., 2023).
+However, these systems are not suitable for NLIDBs as they cannot efficiently handle large relational tables with thousands or millions of rows and diverse data types that are better handled through semantic parsing.
+6. Discussion & Future Work
+DBCopilot has made significant progress in schema routing and schema-agnostic NL2SQL. However, there are some limitations that need to be addressed in future work.
+  1. 1)
+    Cross-Databases Querying.
+While current system shows the capability to route NL questions over massive databases, it is limited to generating SQL queries against a single database.
+Extending DBCopilot to support cross-database SQL querying faces several challenges: i) the heterogeneity of different databases and the complexity of ETL processes for cross-database SQL querying; ii) the increased difficulty in discovering implicit relationships between tables across multiple databases than within a single database; iii) the absence of comprehensive benchmarks for evaluating cross-database NL2SQL systems.
+We envision an alternative direction to decompose complex NL questions into multiple sub-questions, each answerable by a single database, and then integrate the results.
+  2. 2)
+    Synthetic Data Quality.
+As discussed in the case study, the quality of synthetic data used for training has a notable impact on router performance.
+Future efforts could focus on: i) generating high-quality synthetic data by using larger LLMs with better factual consistency; ii) providing more schema descriptions to reduce hallucination; iii) implementing automated validation mechanisms using TNLI (Bussotti et al., 2023) to identify and remove low-quality synthetic data.
+In addition, reducing the amount of data used to train schema routers while maintaining their performance is also a worthwhile exploration.
+  3. 3)
+    Dynamic Schema Update.
+In real-world scenarios, database schemata may evolve over time, which creates a need for all retrieval-based methods to update the index.
+Although we have decoupled the dynamic schema routing into a compact and flexible copilot model, and it currently takes less than 4 hours to train a schema router from scratch on thousands of tables, managing to incremental update DSI can further reduce the computational resources required for evolving schemata.
+7. Conclusion
+In this paper, we present DBCopilot, a framework for effectively scaling natural language querying to massive databases.
+DBCopilot decouples schema-agnostic NL2SQL into schema routing and SQL generation through LLM-Copilot collaboration.
+We present an approach to effectively navigate the target database and tables for an NL question in a relation-aware, end-to-end manner.
+We also propose a training data synthesis paradigm to adapt routers over massive databases without manual intervention, and explore various prompt strategies to incorporate multiple candidate schemata in LLM-based SQL generation.
+Extensive experiments verify the effectiveness of our solution.
+In summary, DBCopilot pushes the boundaries of NL2SQL, enabling researchers to better strategize their quest for data accessibility.
+Acknowledgements.
+We sincerely thank the reviewers for their insightful comments and valuable suggestions.
+This work was supported in part by Beijing Natural Science Foundation under Grant L243006, and in part by the Natural Science Foundation of China under Grant No.
+62476265 and 62306303.